### v2

# Compressible and immiscible fluids with arbitrary density ratios: diffuse interface theory

Abstract
For the water-air system, the bulk density ratio is as high as about 1000; no model can fully tackle such a high density ratio system. In the Navier-Stokes and Euler equations, the density $\rho$ within the water-air interface is assumed to be a constant based on the Boussinesq approximation namely $\rho(\mathrm{d}\mathbf{u}/\mathrm{d}t)$, which does not account for the true momentum evolution $\mathrm{d}(\rho\mathbf{u})/\mathrm{d}t$ ($\mathbf{u}$-fluid velocity). Here, we present an alternative theory for the density evolution equations of immiscible fluids in computational fluid dynamics, differing from the concept of Navier-Stokes and Euler equations. Our derivation is built upon the physical principle of energy minimization from the aspect of thermodynamics. The present results provide a generalization of Bernoulli’s principle for energy conservation and a general formulation for the sound speed. The present model can be applied for immiscible fluids with arbitrarily high density ratios, thereby, opening a new window for computational fluid dynamics both for compressible and incompressible fluids.
I Introduction
Immiscible fluids with a high density ratio are ubiquitous in our daily lives. A typical example is the water-air system, where the density ratio is approximately 1000. However, addressing the density evolution in high-density-ratio systems remains an open question. In the literature, two typical approaches are used to treat density evolution at the water-air interface: (I) the Boussinesq approximation and (II) the interpolation method.
(I) Boussinesq approximation: The first one is the Boussinesq approximation, where the kinetic equation is written as
$\rho\frac{\mathrm{d}\mathbf{u}}{\mathrm{d}t}=-\nabla p+\text{rest terms}.$ (1)
Here, $p$ is the hydrodynamic pressure and $\mathbf{u}$ represents the fluid velocity. In this approach, the density is assumed to be a constant everywhere at a fixed temperature, i.e., $\rho=\text{constant}$. The limitation of the Boussinesq approximation is that the “true” momentum evolution
$\frac{\mathrm{d}(\rho\mathbf{u})}{\mathrm{d}t}$ (2)
has not been taken into account since the density evolution is overlooked. Another limitation of the Boussinesq approximation is the consideration of gravity and buoyancy which are actually caused by the difference in the density. A constant density throughout the system cannot lead to gravity and buoyancy. When gravity and buoyancy are much more pronounced than the inertial force, a density difference term, $\Delta\rho=(\rho_{w}-\rho_{a})\mathbb{g}$ is usually added to “rest terms” in Eq. (1). Here, $\rho_{w}$ and $\rho_{a}$ are the bulk densities of the water and air, respectively, and $\mathbb{g}$ is the gravity acceleration.
(II) Interpolation method: A second approach is the interpolation of the density over the volume concentration $\phi$ of water as Wu et al. (2024); Wörner (2012)
$\rho=\rho_{w}\phi+\rho_{a}(1-\phi).$ (3)
This approach has its advantage of computing gravity and buoyancy implicitly. However, there are two limitations of the interpolation approach: [II(A)]: Conflicts between incompressibility and diffusion. [II(B)]: Uniform excess volume.
  • [II(A)]: Conflicts between incompressibility and diffusion. On the one hand, the incompressible condition requires that
    $\frac{\mathrm{d}\rho}{\mathrm{d}t}=0.$ (4)
    On the other hand, the total time derivative of the density according to the interpolation approach can be written as
    $\frac{\mathrm{d}\rho}{\mathrm{d}t}=(\rho_{w}-\rho_{a})\frac{\mathrm{d}\phi}{\mathrm{d}t}\neq 0.$ (5)
    When diffusion takes place, i.e., $\mathrm{d}\phi/\mathrm{d}t\neq 0$, we have $\mathrm{d}\rho/\mathrm{d}t\neq 0$ if $\rho_{w}\neq\rho_{a}$, so that in the linear interpolation approach, the incompressible condition is not compatible with diffusion for two immiscible phases with distinct densities.
  • [II(B)]: Uniform excess volume. According to the definition of density, we have the following expression for the mixture density at any position $\bm{r}$ in the considered domain $\Omega$ as
    $\rho=\frac{\rho_{w}v_{w}+\rho_{a}v_{a}}{V}=\rho_{w}\frac{v_{w}}{V}+\rho_{a}\frac{v_{a}}{V},$ (6)
    where $V\subset\Omega$ is a volume element including sufficient amount of fluid particles for statistic thermodynamics, and expressed as $V=v_{w}+v_{a}+v_{e}$; $v_{e}$ depicts the excess volume and $v_{i}$ (i=w,ai=w,\ a) is the volume of the $i$ species. By assuming a zero excess volume, i.e., $v_{e}=0$, we have
    $V=v_{w}+v_{a}.$ (7)
    Substituting Eq. (7) into Eq. (6) and using the definition of volume concentration
    $\phi_{i}=v_{i}/V,$ (8)
    The linear interpolation, Eq. (3) is replicated. Suffice it to say, the linear interpolation according to Eq. (3) indicates a zero excess volume of mixing at every position of the considered domain $\Omega$. Any other kinds of interpolations define a certain excess volume, which is also uniform everywhere in the domain $\Omega$. The assumption of uniform excess volume (either zero or non-zero) deviates from the physical fact that the local excess volume $v_{e}$ is affected by the local pressure, which should be space dependent; the excess volume at each position in the domain should have its own characteristics, which will be considered in the current work. A detailed discussion of our concept regarding the excess volume is presented in Section III.
Note that only for the uniform zero excess volume, we have the following constraint
$\phi_{w}+\phi_{a}=1.$ (9)
When $V\neq v_{a}+v_{w}$, the above constraint, Eq. (9) generally does not hold. In summary, to solve the high-density-ratio problem, two evolution equations are required
  • one is for the volume concentration;
  • the other one is for the excess volume or the molar volume of the mixture.
In this work, we explore the high-density ratio problem through the physical principle of energy minimization. We propose a generalized theoretical framework for the evolution equations of fluid dynamics and derive a novel kinetic equation governing density dynamics at the water-air interface. The resulting density evolution equation yields three key contributions: (A): A universal formulation for sound speed; (B): A generalized equation of state for an ideal gas; (C): A generalization of Bernoulli’s law, extending the conservation of potential and kinetic energy. For each contribution, a case study and application is presented in Section VII.
The rest of the paper is structured as follows. In Section II, we provide a brief overview of two high-density ratio models from the literature, clarify the distinction between the total time derivative and the mass conservative derivative, and examine the validity of the conventional continuity equation. Section III outlines the setup of our study and introduces a new concept of volume concentration. In Section IV, we present a dissipation-conservation principle that establishes a connection between mass conservation and energy dissipation. Section V applies this principle to derive the classical fluid dynamics equations for uniform systems. Building on these results, Section VI extends the derivation to non-uniform systems, derives the energy dissipation, and formulates a novel kinetic equation for density. In Section VII, we demonstrate the application of our theory by deriving the sound speed, the equation of state (EOS) for an ideal gas, and the density profile across a water-air interface. Finally, Section VIII concludes the paper with a summary of our findings.
II Two models for high density ratio system in literature
II.1 Ding-Spelt-Shu approach
The Navier-Stokes equation for incompressible fluids reads Ding et al. (2007)
$\rho\frac{\mathrm{d}\mathbf{u}}{\mathrm{d}t}=\rho\bigg(\frac{\partial\mathbf{u}}{\partial t}+\mathbf{u}\cdot\nabla\mathbf{u}\bigg)=-\nabla p+\nabla\cdot[\eta(\nabla\mathbf{u}+\nabla\mathbf{u}^{T})]+\mathbb{f}_{s},$ (10)
where $\eta$ is the dynamic viscosity and $\mathbb{f}_{s}$ depicts the surface tension force. The notation $\nabla\mathbf{u}^{T}$ means $(\nabla\mathbf{u})^{T}$.
The total time derivative of linear momentum is expressed as
$\frac{\mathrm{d}(\rho\mathbf{u})}{\mathrm{d}t}$ $=\frac{\partial(\rho\mathbf{u})}{\partial t}+\nabla\cdot(\rho\mathbf{u}\otimes\mathbf{u})$
$=\frac{\partial\rho}{\partial t}\mathbf{u}+\rho\frac{\partial\mathbf{u}}{\partial t}+(\nabla\cdot\rho\mathbf{u})\mathbf{u}+\rho\mathbf{u}\cdot\nabla\mathbf{u}$
$=\rho\bigg(\frac{\partial\mathbf{u}}{\partial t}+\mathbf{u}\cdot\nabla\mathbf{u}\bigg)+\bigg[\frac{\partial\rho}{\partial t}+\nabla\cdot(\rho\mathbf{u})\bigg]\mathbf{u}.$ (11)
Here, the following vector calculus has been used
$\nabla\cdot(\mathbb{a}\otimes\mathbb{b})=(\nabla\cdot\mathbb{a})\mathbb{b}+\mathbb{a}\cdot\nabla\mathbb{b}.$ (12)
So that an additional time evolution equation for the density is required to iterate the time evolution of the momentum, as shown in Eq. (11).
According to the linear interpolation, as adopted in Ref. Ding et al. (2007), the total time derivative of the density reads
$\frac{\mathrm{d}\rho}{\mathrm{d}t}=(\rho_{w}-\rho_{a})\frac{\mathrm{d}\phi}{\mathrm{d}t}=(\rho_{a}-\rho_{w})\nabla\cdot\mathbf{j},$ (13)
where the diffusion equation with diffusion flux $\mathbf{j}$ has been used
$\frac{\mathrm{d}\phi}{\mathrm{d}t}=\frac{\partial\phi}{\partial t}+\mathbf{u}\cdot\nabla\phi=-\nabla\cdot\mathbf{j}.$ (14)
In Ding-Spelt-Shu approach Ding et al. (2007), a term $(\mathrm{d}\rho/\mathrm{d}t)\mathbf{u}$ is added to both sides of Eq. (10), leading to
$\frac{\mathrm{d}(\rho\mathbf{u})}{\mathrm{d}t}=-\nabla p+\nabla\cdot[\eta(\nabla\mathbf{u}+\nabla\mathbf{u}^{T})]+\mathbb{f}_{s}+(\rho_{a}-\rho_{w})(\nabla\cdot\mathbf{j})\mathbf{u}.$ (15)
When removing $(\mathrm{d}\rho/\mathrm{d}t)\mathbf{u}$ from both sides of Eq. (15), the Navier-Stokes equation, Eq. (10) is replicated.
II.2 Abels-Garcke-Grün approach
The Abels-Garcke-Grün approach Abels et al. (2012) differs from the Ding-Spelt-Shu method. In Abels-Garcke-Grün approach, the momentum equation is assumed to be written as
$\frac{\mathrm{d}(\rho\mathbf{u})}{\mathrm{d}t}=-\nabla p+\nabla\cdot\mathbb{T}+\mathbb{f}_{s},$ (16)
with an unknown stress tensor $\mathbb{T}$. By moving the time derivative of the density to the right hand side, Eq. (16) is rewritten as
$\rho\frac{\mathrm{d}\mathbf{u}}{\mathrm{d}t}=-\nabla p+\nabla\cdot\mathbb{T}+\mathbb{f}_{s}-\frac{\mathrm{d}\rho}{\mathrm{d}t}\mathbf{u}.$ (17)
Substituting Eq. (13) into Eq. (17) leads to
$\rho\frac{\mathrm{d}\mathbf{u}}{\mathrm{d}t}=-\nabla p+\nabla\cdot\mathbb{T}+\mathbb{f}_{s}-(\rho_{a}-\rho_{w})(\nabla\cdot\mathbf{j})\mathbf{u},$ (18)
It is apparent that the last term is not objective that a coordinate transformation, such Newtonian transformation, will change the formulation of the equation Alt (2009). This non-objective problem is solved by Abels, Garcke, and Grün via the following method: By using the vector calculus $\nabla\cdot(\mathbf{j}\otimes\mathbf{u})=(\nabla\cdot\mathbf{j})\mathbf{u}+\mathbf{j}\cdot\nabla\mathbf{u}$, the last term in Eq. (18) is rewritten as
$(\nabla\cdot\mathbf{j})\mathbf{u}=\nabla\cdot(\mathbf{j}\otimes\mathbf{u})-\mathbf{j}\cdot\nabla\mathbf{u}.$ (19)
Abels, Garcke, and Grün define a new stress tensor by combining $\nabla\cdot(\mathbf{j}\otimes\mathbf{u})$ with $\nabla\cdot\mathbb{T}$, namely
$\tilde{\mathbb{T}}=\mathbb{T}+\mathbf{j}\otimes\mathbf{u}:=\eta(\nabla\mathbf{u}+\nabla\mathbf{u}^{T}).$ (20)
In this way, the modified Navier-Stokes equation in the work of Abels-Garcke-Grün is expressed as
$\rho\frac{\mathrm{d}\mathbf{u}}{\mathrm{d}t}=-\nabla p+\nabla\cdot\eta(\nabla\mathbf{u}+\nabla\mathbf{u}^{T})+(\rho_{a}-\rho_{w})\mathbf{j}\cdot\nabla\mathbf{u}.$ (21)
The last term is proportional to the gradient of the velocity, which is now objective. However, the physical meaning for the definition of the new stress tensor $\tilde{\mathbb{T}}$ is unclear. Specially, it is unclear why a symmetric stress tensor $\tilde{\mathbb{T}}$ is obtained by combining a unknown stress tensor $\mathbb{T}$ with an asymmetric stress tensor $\mathbf{j}\otimes\mathbf{u}$ for arbitrary velocity $\mathbf{u}$ and arbitrary diffusion flux $\mathbf{j}$.
Other existing literatures are based either on the continuity equation for the density evolution Tryggvason et al. (2001); Puckett et al. (1997); Fedkiw et al. (1999) (which will be discussed in Section II C) or on the linear interpolation of the density Fakhari et al. (2017); Shukla et al. (2010) according to Eq. (3).
II.3 Total time derivative, mass conservation derivative, and validity of the continuity equation
We often encounter two types of derivatives when deriving the kinetic equation in fluid dynamics: (I) the total time derivative and (II) the conservative derivative. We clarify the difference between these two derivatives, as follows. The total time derivative reads
$\frac{\mathrm{d}\rho}{\mathrm{d}t}=\frac{\partial\rho}{\partial t}+\frac{\partial\rho}{\partial x}\frac{\mathrm{d}x}{\mathrm{d}t}+\frac{\partial\rho}{\partial y}\frac{\mathrm{d}y}{\mathrm{d}t}+\frac{\partial\rho}{\partial z}\frac{\mathrm{d}z}{\mathrm{d}t}=\frac{\partial\rho}{\partial t}+\mathbf{u}\cdot\nabla\rho.$ (22)
The mass conservative derivative reads Tryggvason et al. (2001)
$\frac{\mathrm{D}\rho}{\mathrm{D}t}$ $=\frac{\partial\rho}{\partial t}+\nabla\cdot(\rho\mathbf{u})$
$=\frac{\mathrm{d}\rho}{\mathrm{d}t}+\rho\nabla\cdot\mathbf{u}.$ (23)
When there is no diffusion, i.e., $\mathrm{D}\rho/\mathrm{D}t=0$, we obtain
$\frac{\mathrm{d}\rho}{\mathrm{d}t}=-\rho\nabla\cdot\mathbf{u}.$ (24)
According to Eq. (24), the divergence free of the fluid velocity $\nabla\cdot\mathbf{u}$ is equivalent to the incompressible condition of $\mathrm{d}\rho/\mathrm{d}t=0$, but contradicts with the case with diffusion, where $\mathrm{d}\rho/\mathrm{d}t\neq 0$.
Figure 1: Sketch for the validity of the convection flux in a water-air system. The water-air interface consists of water (blue) and gas (orange) molecules. Within the water-air interface, the density is not uniform. For example, when measuring the convection flux across the reference line (green), the density at the left and right hand side of the reference line is different.
We note that, in general, the so-called mass conservative derivative should not be used to derive the kinetic equation. The key issue is that mass conservation has not been connected to the physical principle of dissipation, as will be demonstrated below. Furthermore, the mass conservative derivative is problematic for deriving the continuity equation, as illustrated in the following. In the absence of diffusion, the convective mass flux $\mathbf{j}_{c}$ is commonly expressed in the literature as
$\mathbf{j}_{c}=\rho\mathbf{u},$ (25)
which has been often adopted to derive the so-called continuity equation
$\frac{\partial\rho}{\partial t}+\nabla\cdot(\rho\mathbf{u})=0.$ (26)
We emphasize that Eq. (25) can only be used to formulate the mass flux caused by convection in the absence of spatial density variations. However, when the density varies in space, this mass flux expression, Eq. (25), becomes invalid. As shown in Fig. 1, we consider a system consisting of water (blue) and gas (orange) molecules. In the bulk of the gas phase, the convection flux across the reference line (highlighted in blue) with a fluid velocity $\mathbf{u}$ can be readily expressed as $\rho\mathbf{u}$. Within the water-air interface, the density varies from $\rho_{1}$ to $\rho_{2}$ from the left to the right side of the reference line (green line). In this case, a unique density $\rho$ cannot be used to measure the convection flux. Suffice it to say, the conventional continuity equation, Eq. (26) cannot be applied in regions where an interface is present. The fundamental problems are two-fold:
  • (P1)
    When there is a spatial variation in density, the relationship between mass conservation and energy dissipation becomes unclear in the continuity equation, Eq. (26).
  • (P2)
    When there is a spatial variation in density, how are the molecules convected from the region of $\rho_{1}$ to the region of $\rho_{2}$, and how do the former one occupy the latter region? Are there any changes of the excess volume as molecules move from the $\rho_{1}$ region to the $\rho_{2}$ region?
III Setup, definition and excess volume concept
We consider an isothermal, closed system composed of two immiscible phases, denoted by $\alpha$ and $\beta$. Each phase contains $K\in\mathbb{Z}$ chemical components within a domain $\Omega$. To illustrate how our formulation differs from existing models, we take a water–gas system as an example. The bulk densities of the water and gas phases are denoted by $\rho_{w}$ and $\rho_{a}$, respectively. We assume that the water–gas interface has a finite thickness, within which water and gas molecules coexist in a mixed region. A representative volume element $V$ is considered, containing a statistically sufficient number of fluid particles in the interfacial region to evaluate the local fluid density as
$\rho=\frac{m_{w}+m_{a}}{V}=\frac{\rho_{w}v_{w}+\rho_{a}v_{a}}{V},$ (27)
where $v_{w}$ and $v_{a}$ denote the partial volumes of the water and gas components, respectively, and $m_{w}$ and $m_{a}$ represent the masses of water and gas contained within the volume element. When $V=v_{w}+v_{a}$, corresponding to an ideal mixture with zero excess volume throughout the system, the overall density $\rho$ can be expressed as a linear interpolation of the volume fractions, defined as $\phi_{i}=v_{i}/V$, as introduced earlier. The assumption of zero excess volume in every representative element across the water–gas interface is commonly adopted in the literature and is often referred to as the incompressible fluid approximation. However, enforcing exactly zero excess volume implies that the density varies monotonically from the bulk water phase to the bulk gas phase. This contradicts recent findings indicating that the density profile across the water–air interface exhibits a non-monotonic behavior Pezzotti et al. (2018); Creazzo and Luber (2024).
Instead of assuming zero excess volume throughout the system, we propose a more general approach (Fig. 2a) in which
$V=v_{w}+v_{a}+v_{e},$ (28)
where $v_{e}\neq 0$ denotes the excess volume, which exhibits spatial variation as a result of local pressure changes. Instead of explicitly solving for $v_{e}$, we redistribute it between $v_{w}$ and $v_{a}$ by an appropriate allocation scheme, resulting in two subregions with volumes $v_{w}^{\prime}$ and $v_{a}^{\prime}$. Consequently, the total volume satisfies $V=v_{w}^{\prime}+v_{a}^{\prime}$ (Fig. 2b). This redistribution of excess volume implies that the local partial densities of water and gas must vary accordingly, since the individual masses of the two components within the volume element $V$ remain constant. We define the volume composition as
$\phi_{w}=\frac{v_{w}^{\prime}}{V}\quad\text{and}\quad\phi_{a}=\frac{v_{a}^{\prime}}{V}.$ (29)
Therefore, the fluid density within the interface is calculated as
$\rho=\rho_{w}^{\prime}\frac{v_{w}^{\prime}}{V}+\rho_{a}^{\prime}\frac{v_{a}^{\prime}}{V}=\rho_{w}^{\prime}\phi+\rho_{a}^{\prime}(1-\phi).$ (30)
The partial densities $\rho_{w}^{\prime}$ and $\rho_{a}^{\prime}$ are implicitly influenced by the excess volume and are therefore no longer constant, as they vary with the redistribution of the excess volume. Consequently, to fully characterize the state of the system, two evolution equations are required: one governing the volume composition $\phi$, and the other governing the density $\rho$.
The classical treatment of incompressible flow corresponds to a special case in which the excess volume vanishes both in the bulk phases and across the interface. Under this assumption, it is sufficient to describe the system using a single variable—either the density $\rho$ or the volume composition $\phi$. A representative example employing $\phi$ is the van der Waals formulation of free energy and its associated diffuse-interface framework Wöhrwag et al. (2018); Ashour et al. (2023); Bestehorn et al. (2021); Jacqmin (1999); Anderson et al. (1998); Yue et al. (2004); Lowengrub and Truskinovsky (1998), in which a single phase order parameter $\phi$ is used to characterize the system’s state and evolution. This well-established approach is commonly referred to as the phase-field model or Cahn–Hilliard model. Another classical example is Euler’s formulation for density evolution, where the standard continuity equation, Eq. (26), is adopted. For incompressible flow, Euler’s formulation reduces to the condition that the convective velocity field is divergence-free. It is important to note that, for incompressible fluids, the van der Waals and Euler formulations cannot be applied simultaneously, as doing so would lead to an overdetermined system.
Figure 2: The volume concentration within the volume element $V$. (a) Conventional definition. The volume element $V$ is divided into the volume of air ($v_{a}$) and the volume of water ($v_{w}$), plus the excess volume $v_{e}$. In this definition, the densities $\rho_{a}$ and $\rho_{w}$ are constants. (b) Current concept. The volume element $V$ is divided into the volume of air ($v_{a}^{\prime}$) and the volume of water ($v_{w}^{\prime}$). We have distributed the excess volume into the volumes of $v_{a}^{\prime}$ and $v_{w}^{\prime}$. By this way, the densities $\rho_{a}^{\prime}$ and $\rho_{w}^{\prime}$ are affected by the local excess volume and no more constants.
The present work develops a generalized theoretical framework for systems with nonzero excess volume within the fluid–fluid interface, particularly for cases involving large density ratios. The objective is to extend both the classical van der Waals and Euler formulations by incorporating interfacial compressibility. Here, compressibility refers to the variation of density across the interface arising from mixing with a finite excess volume. Even when the excess volume is nonzero, if $\rho_{w}^{\prime}\approx\rho_{a}^{\prime}=\rho^{*}$, the single-variable van der Waals formulation remains adequate to represent the compressibility, since $\rho\approx\rho^{*}$ according to Eq. (30). Either the variation of $\rho^{*}$ or $\phi$ can then describe the compressibility. However, when $\rho_{w}^{\prime}\gg\rho_{a}^{\prime}$, the van der Waals formulation loses its validity and must be generalized to properly describe the evolution of the density field.
The above concept can be readily extended to the more general case of a multicomponent system, for which the following equality holds:
$\phi_{i}=\frac{v_{i}^{\prime}}{V}\quad\text{and}\quad\sum_{i}v_{i}^{\prime}=V,\quad\text{and}\quad\sum_{i=1}^{K}\phi_{i}=1.$ (31)
The boundary of the domain $\Omega$ is denoted by $\Gamma$. For a binary system, we use the notation $\phi_{1}=\phi$, $\phi_{2}=1-\phi$. The density of the mixture at any position $\bm{r}\in\Omega$ is denoted by $\rho(\bm{r})$.
IV Dissipation principle
In this section, we present a dissipation-conservation theorem, which will be used to derive the kinetic equation in the following sections.
The dissipation-conservation theorem describes the energy dissipation for the field $u$ and the associated potential $\Phi$. It establishes a connection between the first law of thermodynamics (conservation) with the second law of thermodynamics (dissipation). The energy is formulated as a function of $u$ as
$E=\int_{\Omega}e(u)\mathrm{d}\Omega.$ (32)
The time derivative reads
$\frac{\mathrm{d}E}{\mathrm{d}t}=\int_{\Omega}\frac{\mathrm{d}e}{\mathrm{d}u}\frac{\mathrm{d}u}{\mathrm{d}t}\mathrm{d}\Omega=\int_{\Omega}\Phi\frac{\mathrm{d}u}{\mathrm{d}t}\mathrm{d}\Omega,$ (33)
where we have defined a potential $\Phi=\mathrm{d}e/\mathrm{d}u$. From Eq. (33), we obtain the conserved evolution of $u$ as
$\frac{\mathrm{d}u}{\mathrm{d}t}=\nabla\cdot\tau_{0}\nabla\Phi=\nabla\cdot\tau_{0}\nabla\frac{\mathrm{d}e}{\mathrm{d}u},$ (34)
where $\tau_{0}$ is a positive constant. Substituting Eq. (34) into Eq. (33), integrating by parts, and using the no-flux boundary condition, we obtain
$\frac{\mathrm{d}E}{\mathrm{d}t}=\int_{\Omega}\frac{\mathrm{d}e}{\mathrm{d}u}\frac{\mathrm{d}u}{\mathrm{d}t}\mathrm{d}\Omega=\int_{\Omega}\frac{\mathrm{d}e}{\mathrm{d}u}\nabla\cdot\tau_{0}\nabla\Phi\mathrm{d}\Omega=-\int_{\Omega}\tau_{0}(\nabla\Phi)^{2}\mathrm{d}\Omega\leq 0.$ (35)
In summary, an energy dissipation in the form of Eq. (33) simultaneously implies a conservation equation in the form of Eq. (34), and vice versa.
From Eq. (34) and Eq. (35), we observe that the system energy is non-increasing as long as $\tau_{0}\geq 0$. The physical interpretation of $\tau_{0}$ can be understood as follows. Applying the chain rule, we rewrite Eq. (34) in the form of a diffusion equation:
$\frac{\mathrm{d}u}{\mathrm{d}t}=\nabla\cdot\bigg(\tau_{0}\frac{\mathrm{d}^{2}e}{\mathrm{d}u^{2}}\nabla u\bigg)=\nabla\cdot(\mathcal{D}\nabla u).$ (36)
Thus, the coefficient $\tau_{0}$ establishes a connection between the purely dynamic parameter $\mathcal{D}$ (independent of conservative energy) and the energy-dependent term $\frac{\mathrm{d}^{2}e}{\mathrm{d}u^{2}}$ via the relation $\tau_{0}=\mathcal{D}/\frac{\mathrm{d}^{2}e}{\mathrm{d}u^{2}}$. In other words, $\tau_{0}$ represents the ratio of non-conservative to conservative energy contributions to the kinetics of the system.
The dissipation equation for non-conserved variables can be written in the steepest descendent approach:
$\frac{\mathrm{d}u}{\mathrm{d}t}=-\tau_{0}u.$ (37)
V Uniform system: Derivation of Euler’s equation and Newton’s law
In this section, we apply the dissipation-conservation principle to a uniform system to derive the conventional fluid dynamics equations. We suppose that the total energy of the system consists of the free energy $\mathcal{F}$, the pressure energy $\mathcal{P}$, and the kinetic energy $\mathcal{K}$
$\mathcal{E}=\mathcal{F}+\mathcal{P}+\mathcal{K}.$ (38)
The free energy, pressure energy, and kinetic energy are integrated over the energy density as
$\mathcal{F}$ $=\int_{\Omega}f(\phi,T)\mathrm{d}\Omega,$ (39)
$\mathcal{P}$ $=\int_{\Omega}p\mathrm{d}\Omega,$ (40)
$\mathcal{K}$ $=\int_{\Omega}k\mathrm{d}\Omega=\int_{\Omega}\rho\mathbf{u}\cdot\mathbf{u}\mathrm{d}\Omega.$ (41)
The free energy density depends on the volume concentration $\phi$ and temperature $T$, and may be expressed as $f=\theta-Ts$, where $\theta$ denotes the internal energy and $s\sim\phi\ln\phi$ represents the mixing entropy, which is distinct from the thermal (heat) entropy.
The total time derivative of the total energy reads
$\frac{\mathrm{d}\mathcal{E}}{\mathrm{d}t}=\int_{\Omega}\frac{\partial f}{\partial T}\frac{\mathrm{d}T}{\mathrm{d}t}+\frac{\partial f}{\partial\phi}\frac{\mathrm{d}\phi}{\mathrm{d}t}+\frac{\mathrm{d}p}{\mathrm{d}t}+\frac{\mathrm{d}(\rho\mathbf{u})}{\mathrm{d}t}\cdot\mathbf{u}+\rho\mathbf{u}\cdot\frac{\mathrm{d}\mathbf{u}}{\mathrm{d}t}\mathrm{d}\Omega.$ (42)
Next, by using the chain rule for the space dependent pressure $p(\mathbf{r})$, we rewrite the total time derivative of the pressure as
$\frac{\mathrm{d}p}{\mathrm{d}t}=\frac{\mathrm{d}p}{\mathrm{d}\mathbf{r}}\cdot\frac{\mathrm{d}\mathbf{r}}{\mathrm{d}t}=\nabla p\cdot\mathbf{u}.$ (43)
Substituting Eq. (43) into Eq. (42), we obtain
$\frac{\mathrm{d}\mathcal{E}}{\mathrm{d}t}=\int_{\Omega}\frac{\partial f}{\partial T}\frac{\mathrm{d}T}{\mathrm{d}t}+\frac{\partial f}{\partial\phi}\frac{\mathrm{d}\phi}{\mathrm{d}t}+\bigg[\nabla p+\frac{\mathrm{d}(\rho\mathbf{u})}{\mathrm{d}t}\bigg]\cdot\mathbf{u}+\rho\mathbf{u}\cdot\frac{\mathrm{d}\mathbf{u}}{\mathrm{d}t}\mathrm{d}\Omega.$ (44)
Applying the dissipation principle stated in Section IV to the first part of Eq. (44), we obtain Fourier’s heat equation for the temperature as
$\frac{\mathrm{d}T}{\mathrm{d}t}=\nabla\cdot M_{T}\nabla\frac{\partial f}{\partial T}=\nabla\cdot M_{T}(\partial^{2}f/\partial T^{2})\nabla T=\nabla\cdot D_{T}\nabla T,$ (45)
where the thermal diffusivity $D_{T}$ is defined in terms of the mobility $M_{T}$ as $D_{T}=M_{T}(\partial^{2}f/\partial T^{2})$. Since the thermal diffusivity generally exceeds the mass diffusivity by at least three orders of magnitude, the temperature evolution equation is consequently neglected in the following derivation. For instance, at room temperature, the thermal diffusivity of water is approximately $0.14\times 10^{-6}$ m2/s, whereas its mass diffusivity is only $2.3\times 10^{-9}$ m2/s.
Applying the dissipation principle to the second part of Eq. (44), we obtain a conserved time evolution equation for $\phi$, which is the Fick’s diffusion equation, as
$\frac{\mathrm{d}\phi}{\mathrm{d}t}=\nabla\cdot M\nabla\frac{\partial f}{\partial\phi}=\nabla\cdot M(\partial^{2}f/\partial\phi^{2})\nabla\phi=\nabla\cdot D\nabla\phi,$ (46)
where the mass diffusivity $D$ is defined in terms of the mobility $M$ as $D=M(\mathrm{d}^{2}f/\mathrm{d}\phi^{2})$.
When considering that there is no velocity dissipation, e.g., inviscid fluids, we obtain the momentum balance equation according to the second term in Eq. (44)
$\frac{\mathrm{d}(\rho\mathbf{u})}{\mathrm{d}t}=-\nabla p,$ (47)
which depicts the exchange of the kinetic energy with the potential energy via the work done by the pressure $\nabla p\cdot\mathbf{u}$ and the work done by the inertial force, $\mathbf{f}_{\rho}\cdot\mathbf{u}=\rho\mathbf{u}\cdot\mathbf{u}$.
When the density is assumed to remain constant over time,
$\frac{\mathrm{d}\rho}{\mathrm{d}t}=0,$ (48)
the momentum balance equation is written as
$\rho\frac{\mathrm{d}\mathbf{u}}{\mathrm{d}t}=-\nabla p,$ (49)
which is nothing but the Euler’s equation in fluid dynamics. An integration over the whole domain, we obtain the well-known Newton’s law
$\mathbf{F}=m\mathbf{a},$ (50)
where the mass $m$, the acceleration $\mathbf{a}$, and the force $\mathbf{F}$ are defined as
$m=\int_{V}\rho\mathrm{d}V,\ \mathbf{a}=\frac{\mathrm{d}\mathbf{u}}{\mathrm{d}t},\ \mathbf{F}=\int_{V}-\nabla p\mathrm{d}V.$ (51)
VI Non-uniform system: System energy with an interface
In this section, we apply the dissipation-conservation principle to non-uniform systems and derive the corresponding kinetic equations.
VI.1 Energy formulation and dissipation
We propose that the total energy of the system consists of free energy $\mathcal{F}$, pressure energy $\mathcal{P}$, macroscopic kinetic energy $\mathcal{K}$, and wall free energy $\mathcal{W}$, reading
$\mathcal{E}=\mathcal{F}+\mathcal{P}+\mathcal{K}+\mathcal{W}.$ (52)
In contrast to the uniform system, the free energy of a non-uniform system depends not only on $\phi$ but also on $\nabla\phi$, as derived by Cahn Cahn and Hilliard (1958) and van der Waals Rowlinson (1979), known as van der Waals-Cahn theory. The integration over the local free energy density $e(\bm{\phi},\nabla\bm{\phi})$ leads to
$\mathcal{F}(\bm{\phi},\nabla\bm{\phi})=\int_{\Omega}e(\bm{\phi},\nabla\bm{\phi})d\Omega,$ (53)
$e(\bm{\phi},\nabla\bm{\phi})=f(\bm{\phi})+\sum_{i=1}^{K}\frac{1}{2}\kappa_{i}(\nabla\phi_{i})^{2},$ (54)
where $f(\bm{\phi})$ depicts the bulk free energy density and $\kappa_{i}$ represents the gradient energy coefficient which is related to the interfacial tension of the $\alpha$-$\beta$ interface, $\sigma_{\alpha\beta}$.
Figure 3: Origin of the pressure energy for uniform and nonuniform systems. For a uniform system (a), the pressure at any position is the same, which is taken as a reference value. In a non-uniform system with density variation (b), the pressure is generally non-uniform due to the asymmetric collisions of different sized molecules. Blue and red spheres depict two molecules or atoms with distinct sizes, leading to a variation of the density in space. The arrows sketch the thermal motion of the molecules/atoms.
The $\mathcal{P}=\int_{\Omega}pd\Omega$ term, often overlooked previously and introduced here, describes the pressure energy. We formulate the pressure energy density as Chen and Doolen (1998)
$-p=e-\sum_{i=1}^{K}\mu_{i}\phi_{i},$ (55)
where $\mu_{i}$ is the chemical potential of species $i$, the formulation of which will be derived in the following via the energy dissipation principle. Herein, an essential consideration is that the pressure $p$ is always associated with the free energy $\mathcal{F}$. A physical interpretation of the pressure is shown in Fig. 3. A homogeneous mixture of two distinct species (Fig. 3a) leads to a uniform pressure $p_{0}(\bm{r})=\text{const},\ \forall\bm{r}\in\Omega$, which can be taken as a reference value. However, when the two species are not homogeneously mixed (Fig. 3b), the intrinsic pressure is non-uniform in space; the non-uniform pressure is caused, for example, by the collisions of different molecules with distinct sizes due to the asymmetry in space. In the special case of a homogeneous mixture, the pressure defined in Eq. (55) is consistent with the Landau potential (sometimes called the thermodynamic pressure Wang et al. (2023)), given by
$\Psi=f(\bm{\phi})-\sum_{i=1}^{K}(\partial f/\partial\phi_{i})\phi_{i}.$ (56)
The $\mathcal{K}$ term denotes the kinetic energy of the fluid, which is formulated as
$\mathcal{K}=\int_{\Omega}\bm{\zeta}\cdot\mathbf{u}d\Omega,$ (57)
where $\bm{\zeta}$ is the linear momentum density expressed as $\bm{\zeta}=\rho\mathbf{u}$.
The $\mathcal{W}$ term in Eq. (52) represents the wall free energy, which is integrated over the wall free energy density $f_{w}$ on the boundary $\Gamma$ as $\mathcal{W}=\int_{\Gamma}f_{w}d\Gamma$. The formulation of the wall free energy density $f_{w}$ has been comprehensively discussed elsewhere Wang and Nestler (2024); Wang et al. (2023); Wang and Nestler (2021) and is not the scope of this work.
In the following, we derive the equilibrium condition and the kinetic equations based on the principle of energy minimization, namely
$\frac{\mathrm{d}\mathcal{E}}{\mathrm{d}t}=\dot{\mathcal{E}}\leq 0.$
Next, we evaluate the total time derivative of the free energy $\mathcal{F}$, the pressure energy $\mathcal{P}$, and the kinetic energy $\mathcal{K}$. By using the chain rule and noting that the functional $\mathcal{F}$ has two arguments, $\bm{\phi}$ and $\nabla\bm{\phi}$, the total time derivative of the potential free energy $\mathcal{F}$ reads
$\dot{\mathcal{F}}(\bm{\phi},\nabla\bm{\phi})=\int_{\Omega}\sum_{i=1}^{K}\frac{\partial e}{\partial\phi_{i}}\frac{\mathrm{d}\phi_{i}}{\mathrm{d}t}+\frac{\partial e}{\partial\nabla\phi_{i}}\cdot\frac{\mathrm{d}\nabla\phi_{i}}{\mathrm{d}t}d\Omega.$ (58)
By using the vector calculus $\nabla\cdot(\nabla\phi_{i}\otimes\nabla\phi_{i})=\nabla^{2}\phi_{i}\nabla\phi_{i}+(\nabla\phi_{i}\cdot\nabla)\nabla\phi_{i}$, the definition of materials derivative $\dot{\nabla\phi_{i}}=\partial_{t}\nabla\phi_{i}+\mathbf{u}\cdot\nabla\nabla\phi_{i}$, and integration by parts with no-flux boundary condition for a closed system (more details can be found in Wang et al. (2023); Zhang et al. (2024)), we obtain the following equations
$\dot{\mathcal{F}}=\int_{\Omega}\sum_{i=1}^{K}\mu_{i}\dot{\phi}_{i}+(\nabla\cdot\underline{\underline{\bm{\Theta}}})\cdot\mathbf{u}d\Omega;$ (59)
$\mu_{i}=\frac{\partial e}{\partial\phi_{i}}-\nabla\cdot\frac{\partial e}{\partial\nabla\phi_{i}};$ (60)
$\underline{\underline{\bm{\Theta}}}=\sum_{i}\frac{\partial e}{\partial\nabla\phi_{i}}\otimes\nabla\phi_{i}.$ (61)
The first part of Eq. (59) defines a generalized formulation for the chemical potential, as stated in Eq. (60). At thermodynamic equilibrium, the chemical potential is a constant value throughout the system. In a special case of a zero chemical potential, i.e., $\mu_{i}=0$, Eq. (60) is equivalent to the Euler-Lagrange equation and replicates the Young-Laplace equation; Eq. (60) also provides a physical interpretation for the variational derivative in mathematics. The second part of Eq. (59) defines a stress tensor $\underline{\underline{\bm{\Theta}}}$, which is responsible for the transformation of the potential free energy into the macroscopic kinetic energy Jacqmin (1999). This stress tensor is consistent with the Korteweg stress by evaluating the derivative via considering Eq. (54), namely $\partial_{\nabla\phi_{i}}e=\kappa_{i}\nabla\phi_{i}$.
Noteworthily, a uniform chemical potential throughout the system cannot well define the thermodynamic equilibrium; the pressure evolution should be considered as well. An evaluation of the time derivative for the pressure energy and the kinetic energy leads to
$\dot{\mathcal{P}}=\int_{\Omega}\frac{\mathrm{d}p(\mathbf{r})}{\mathrm{d}t}\mathrm{d}\Omega=\int_{\Omega}\nabla p\cdot\dot{\mathbf{r}}\mathrm{d}\Omega=\int_{\Omega}\nabla\cdot(p\mathbf{I})\cdot\mathbf{u}\mathrm{d}\Omega,$ (62)
$\dot{\mathcal{K}}=\int_{\Omega}(\dot{\bm{\zeta}}\cdot\mathbf{u}+\bm{\zeta}\cdot\dot{\mathbf{u}})\mathrm{d}\Omega,$ (63)
where $\mathbf{I}$ is an identity tensor. For the derivation of Eq. (62) and Eq. (63), we have used the chain rule and the definition of the velocity, $\dot{\mathbf{r}}=\mathbf{u}$. Summarizing Eq. (59), Eq. (60), Eq. (61), Eq. (62), and Eq. (63) and collecting all the $\cdot\mathbf{u}$ terms, we obtain the total time derivative of the system energy $\mathcal{E}$ as
$\dot{\mathcal{E}}=\int_{\Omega}\sum_{i=1}^{K}\underbrace{\mu_{i}\dot{\phi_{i}}}_{\text{I}}+\underbrace{[\nabla\cdot(\underline{\underline{\bm{\Theta}}}+p\mathbf{I})+\dot{\bm{\zeta}}]}_{\text{II}}\cdot\mathbf{u}+\underbrace{\bm{\zeta}\cdot\dot{\mathbf{u}}}_{\text{III}}d\Omega.$ (64)
The physical interpretation for the three parts of the energy evolution in Eq. (64) is sketched in Fig. 4.
Figure 4: The energy dissipation mechanisms in Eq. (64). I: The diffusion via the exchange of the molecules and vacancies leads to the dissipation of the free energy $\mathcal{F}$. II: The gradient of the free energy $\mathcal{F}$ and the associated pressure $\mathcal{P}$ result in an increase in the momentum $\bm{\zeta}=\rho\mathbf{u}$. III: The difference in the velocity, density, and viscosity in distinct subdomains (dashed rectangle) gives rise to the dissipation of the velocity $\mathbf{u}$ caused by the ‘spring’ drag effect; the spring symbol denote the interaction between the atoms/molecules in different subdomains.
VI.2 Kinetic equations
According to the dissipation-conservation theorem, the first part in Eq. (64), I, defines the generalized diffusion-convection equation in a non-uniform system as
$\dot{\phi_{i}}$ $=\frac{\mathrm{d}\phi_{i}}{\mathrm{d}t}=\frac{\partial\phi_{i}}{\partial t}+\mathbf{u}\cdot\nabla\phi_{i}=\nabla\cdot\varsigma_{i}\nabla\mu_{i}.$ (65)
The coefficient $\varsigma_{i}$ is the mobility.
The second part in Eq. (64), II, reveals the momentum conservation
$\dot{\bm{\zeta}}=\frac{\mathrm{d}(\rho\mathbf{u})}{\mathrm{d}t}=-\nabla\cdot(\underline{\underline{\bm{\Theta}}}+p\mathbf{I}).$ (66)
The conservative force at the right hand side is responsible for the exchange of the potential energy $\mathcal{F}+\mathcal{P}$ with the kinetic energy $\mathcal{K}$. This result is nothing but Newton’s law of momentum balance. Here, a more clear physical meaning of the force is presented; the force term consists of the pressure $p$ and the stress tensor $\underline{\underline{\bm{\Theta}}}$. The stress tensor is due to the gradient term in the potential energy $\mathcal{F}$, while the pressure is contributed by the associated free energy. Note that a viscous dissipation term cannot be directly added to the momentum conservation equation, as there is no quantity corresponding to “viscous energy”.
The only term left in the time evolution of system energy, III, is $\int_{\Omega}\bm{\zeta}\cdot\dot{\mathbf{u}}\mathrm{d}\Omega$. By applying the same dissipation principle as done for the scalar field in Eq. (65), we obtain the dissipation equation for the velocity as
$\dot{\mathbf{u}}=\nabla\cdot\tau_{1}(\nabla\bm{\zeta}+\nabla\bm{\zeta}^{T})+\nabla\cdot\tau_{2}(\nabla\cdot\bm{\zeta})\mathbf{I}.$ (67)
Unlike the dissipation rule for the scalar variable $\phi_{i}$ with the associated potential $\nabla\mu_{i}$ (see Eq. (65)), three terms $\nabla\bm{\zeta}$, $\nabla\bm{\zeta}^{T}$, and $(\nabla\cdot\bm{\zeta})\mathbf{I}$ have to be taken into account for the dissipation of a vector $\mathbf{u}$.
The right hand side of Eq. (67) is equivalent to Newton viscous stress tensor if we define the mobilities $\tau_{1}$ and $\tau_{2}$ as
$\tau_{1}=\eta/\rho^{2},\ \tau_{2}=\nu/\rho^{2},$ (68)
where $\eta$ and $\nu$ are the shear and bulk viscosities Tanaka (2000), respectively.
The dissipation equation, Eq. (67), is essentially a diffusion equation applied to the vector $\mathbf{u}$, as evident from the following comparison:
$\nabla\cdot\tau_{1}\nabla\bm{\zeta}=\nabla\cdot\tau_{1}\nabla(\partial\mathcal{K}/\partial\mathbf{u});\ \bm{\zeta}=\partial\mathcal{K}/\partial\mathbf{u};$ (69)
$\nabla\cdot\varsigma\nabla\mu=\nabla\cdot\varsigma\nabla(\partial\mathcal{F}/\partial\mathbf{\phi});\ \mu=\partial\mathcal{F}/\partial\phi.$ (70)
The linear momentum, $\bm{\zeta}$, plays a role analogous to that of the chemical potential, $\mu$.
A further question is that if the velocity is a conserved variable or not. If not, the dissipation has to be modified by the non-conserved form as
$\frac{\mathrm{d}\mathbf{u}}{\mathrm{d}t}=-\tau_{0}\bm{\zeta},$ (71)
where $\tau_{0}$ is a positive mobility. The non-conserved form is known as the Allen-Cahn type dissipation following the gradient descent path, being consistent with the Langevin equation.
VI.3 The problem of the conventional kinetic energy formulation
Conventional formulation for the kinetic energy is $\tilde{\mathcal{K}}=\int_{\Omega}\frac{1}{2}\rho\mathbf{u}^{2}\mathrm{d}\Omega$. In this case, the time derivative of the kinetic energy reads
$\frac{\mathrm{d}{\tilde{\mathcal{K}}}}{\mathrm{d}t}=\int_{\Omega}\bigg(\frac{1}{2}\frac{\mathrm{d}\rho}{\mathrm{d}t}\mathbf{u}^{2}+\rho\mathbf{u}\cdot\frac{\mathrm{d}\mathbf{u}}{\mathrm{d}t}\bigg)\mathrm{d}\Omega.$ (72)
A noteworthy shortcoming of Eq. (72) is the missing information of the momentum balance as well as the transformation between the potential and kinetic energies. Moreover, the first part in Eq. (72) requires a prior knowledge about the time evolution of the density, which is unknown for two immiscible phases. The density evolution can be avoided if the kinetic energy is written as $\rho\mathbf{u}\cdot\mathbf{u}=\bm{\zeta}\cdot\mathbf{u}$; in this case, the time evolution of the momentum $\bm{\zeta}$ as a whole term is sufficient, subjected to the momentum conservation. Noteworthily, the so-called continuity equation $\partial_{t}\rho+\nabla\cdot(\rho\mathbf{u})=0$ cannot be applied here for density evolution. The reason is as follows: (I) The classic continuity equation is derived for pure fluid, where there is no density variation in space. (II) When considering two immiscible fluids, the classic continuity equation cannot be straightforwardly adopted, since the density varies across the interface. In this case, the mass conservation and energy dissipation are not independent. Suffice it to say, one cannot directly apply a mass conservation without considering the energy dissipation across the interface. (III) Within the interface of two immiscible fluids, one cannot use the condition $\nabla\cdot\mathbf{u}=0$ which is used for pure fluids. The divergence free condition $\nabla\cdot\mathbf{u}=0$ implies a uniform excess volume everywhere in the whole domain $\Omega$ and that the mass is uniformly distributed in the speed of sound. In other words, when the fluid velocity is much less than the speed of sound, the density almost has no variation in space; only in this case, the condition $\nabla\cdot\mathbf{u}=0$ can be applied. For pure homogeneous fluids, the condition $\nabla\cdot\mathbf{u}=0$ is used to solve the unknown pressure $p_{0}$, which is homogeneous and can be considered as a reference value (Fig. 3). However, for two immiscible fluids, the pressure is inhomogeneous across the interface, which cannot be solved simply by $\nabla\cdot\mathbf{u}=0$ due to the density variation.
VI.4 Density evolution equation and the generalization of Bernoulli’s law
Eq. (66) subtracting Eq. (67) with the relation $\dot{\bm{\zeta}}=\dot{\rho}\mathbf{u}+\rho\dot{\mathbf{u}}$, we obtain the temporal equation for the density as
$\mathbf{u}\dot{\rho}=-\nabla\cdot[(p\mathbf{I}+\underline{\underline{\bm{\Theta}}})]-\rho\nabla\cdot[\tau_{1}(\nabla\bm{\zeta}+\nabla\bm{\zeta}^{T})+\tau_{2}(\nabla\cdot\bm{\zeta})\mathbf{I}].$ (73)
As demonstrated in the following, Eq. (73) provides a generalization of Bernoulli’s law.
Considering inviscid fluids via setting $\tau_{1}=\tau_{2}\approx 0$, and the steady state, $\dot{\rho}=\mathbf{u}\cdot\nabla\rho$ where we set $\partial_{t}\rho=0$, we rewrite Eq. (73) as
$\mathbf{u}^{2}\nabla\rho=-[\nabla\cdot(p\mathbf{I}+\underline{\underline{\bm{\Theta}}})].$ (74)
By considering a one-dimensional setup and using the relation $\mathrm{d}(\rho u^{2})=u^{2}\mathrm{d}\rho+2\rho udu$, Eq. (74) is further simplified as
$\mathrm{d}(p+\rho u^{2})=2u\mathrm{d}\rho.$ (75)
This result is nothing but a generalized Bernoulli’s equation. When there is no density variation, namely, $\mathrm{d}(p+\rho u^{2})=0$, we have
$p+\rho u^{2}=\text{constant},$ (76)
which is Bernoulli’s principle. However, when density varies, $\mathrm{d}\rho\neq 0$, the associated energy is changed as well; in this case, the classic Bernoulli’s equation loses its validity.
VI.5 Replication of the Navier-Stokes equations
In this section, we show that in a special case, the present theory is consisent with Navier-Stokes equations. Only in this section, we will consider the factor $1/2$ for the kinetic energy $(1/2)\rho\mathbf{u}\cdot\mathbf{u}$.
The momentum balance equation with the factor $1/2$ reads
$\frac{1}{2}\frac{\mathrm{d}(\rho\mathbf{u})}{\mathrm{d}t}=\frac{1}{2}\rho\frac{\mathrm{d}\mathbf{u}}{\mathrm{d}t}+\frac{1}{2}\mathbf{u}\frac{\mathrm{d}\rho}{\mathrm{d}t}=-\nabla p.$ (77)
For incompressible fluids $\mathrm{d}\rho/\mathrm{d}t=0$, Eq. (77) reduces to
$\frac{1}{2}\rho\frac{\mathrm{d}\mathbf{u}}{\mathrm{d}t}=-\nabla p.$ (78)
Multiplying both sides of the dissipation equation, Eq. (67) by $\rho$, the dissipation equation reads [a factor $1/2$ has been added due to the kinetic energy formulation, $(1/2)\rho\mathbf{u}\cdot\mathbf{u}$ ]
$\frac{1}{2}\rho\frac{\mathrm{d}\mathbf{u}}{\mathrm{d}t}=\rho\{\nabla\cdot\tau_{1}[\nabla(\rho\mathbf{u})+\nabla(\rho\mathbf{u})^{T}]+\nabla\cdot\tau_{2}\nabla\cdot(\rho\mathbf{u})\mathbf{I}\}=\nabla\cdot\eta(\nabla\mathbf{u}+\nabla\mathbf{u}^{T})+\nabla\cdot\nu\nabla\cdot\mathbf{u})\mathbf{I}.$ (79)
Summing Eq. (78) with Eq. (79), we obtain the classic Navier-Stokes equation for incompressible fluids as
$\rho\frac{\mathrm{d}\mathbf{u}}{\mathrm{d}t}=-\nabla p+\nabla\cdot\eta(\nabla\mathbf{u}+\nabla\mathbf{u}^{T})+\nabla\cdot\nu(\nabla\cdot\mathbf{u})\mathbf{I}.$ (80)
We stress that viscous dissipation inherently contradicts the incompressibility condition. When viscous dissipation occurs, density cannot remain constant, as dissipative processes—such as heat release, light emission, or electrical discharge—introduce entropy changes that modify the density. However, under dissipation-free conditions (e.g., isentropic flow), density can reasonably be treated as constant. In this way, the Navier-Stokes equation has the conflicts between viscous dissipation and invariance of density (incompressibility). In fact, Eq. (78) cannot be summed with Eq. (79) because:
  • Eq. (78) is based on the incompressible condition that the density does not vary with time;
  • Eq. (79) depicts a kind of viscous dissipation due to entropy increase. Strictly speaking, when viscous dissipation occurs, the density cannot be a constant.
When there is viscous dissipation, we must use Eq. (77) with the total time derivative of the momentum, rather than Eq. (78). To solve the total time derivative of the momentum, a key difficulty here is that we need prior knowledge about the total time derivative of the velocity as well as the total time derivative of the density. The former one is related to the velocity dissipation. Once the dissipation formulation is known, the time evolution of the density can be deduced.
The dissipation formulation, Eq. (79) is particularly chosen to be consistent with the Newton viscous stress tensor to dissipate the velocity and the associated energy as
$\int_{\Omega}\rho\mathbf{u}\cdot\frac{\mathrm{d}\mathbf{u}}{\mathrm{d}t}\mathrm{d}\Omega=-\tau_{1}\int_{\Omega}\nabla(\rho\mathbf{u}):\nabla(\rho\mathbf{u})\mathrm{d}\Omega-\tau_{2}\int_{\Omega}[\nabla\cdot(\rho\mathbf{u})]^{2}\mathrm{d}\Omega\leq 0.$ (81)
Other dissipation formations for $\mathrm{d}\mathbf{u}/\mathrm{d}t$ are also possible as long as the integration
$\int_{\Omega}\rho\mathbf{u}\cdot\frac{\mathrm{d}\mathbf{u}}{\mathrm{d}t}\mathrm{d}\Omega$ (82)
is not greater than 0.
VI.6 Remarks
The key point of the present theory is the energy minimization based on an alternative formulation for the kinetic energy, $\rho\mathbf{u}\cdot\mathbf{u}$, rather than the conventional formulation, $\frac{1}{2}\rho\mathbf{u}^{2}$. In our derivation, the dissipation of the kinetic energy is divided into two parts. The first part is the momentum conservation, $\mathrm{d}(\rho\mathbf{u})/\mathrm{d}t$ which changes not only the fluid velocity $\mathbf{u}$ but also the density $\rho$. This derivation differs from Euler and Navier-Stokes equations where the momentum variation is only caused by the evolution of the fluid velocity. The second part of the kinetic energy dissipation is the velocity dissipation due to the difference in the density, viscosity, and velocity, as explained in Fig. 4III.
Another vital point of the current work is the pressure energy, the formulation of which is consistent with many pioneers Chen and Doolen (1998). The physical meaning of the pressure has been elucidated in Fig. 3. The minimization of the pressure energy leads to the so-called grand potential model Wang et al. (2012). In the lattice model, the free energy and the pressure can be explicitly formulated according to the theory of thermodynamics, such as Landau potential. However, for non-lattice models, the pressure cannot be explicitly identified. This case is the consideration in Euler and Navier-Stokes equations, where the unknown pressure is solved by the divergence free of the velocity for incompressible flow. For liquid and solid phases, the lattice model can be applied because of the regularized order of the structure within the short range. For gas phases, the lattice model loses its validity due to the disorder of the structure; one has to use the equation of state to formulate the partial pressure of the gas phase Wöhrwag et al. (2018); Ledesma-Aguilar et al. (2014).
In the current work, we consider the linear momentum $\bm{\zeta}=\rho\mathbf{u}$. Non-linear terms related to fluid velocity may be accounted for the dissipation if we consider non-linear momentum, for example, $\bm{\zeta}\propto\mathbf{u}^{n}$ or $\bm{\zeta}\propto(\nabla\mathbf{u})^{n}$. The non-linear momentum not only modifies the velocity dissipation but also changes the momentum balance, which may be applied for understanding highly nonlinear effects, such as turbulence. These specialized cases are out of the scope of the present discussion.
Let us recall the momentum balance equation
$\rho\frac{\mathrm{d}\mathbf{u}}{\mathrm{d}t}+\mathbf{u}\frac{\mathrm{d}\rho}{\mathrm{d}t}=-\nabla p.$ (83)
The evolution of the velocity corresponds to the dissipation according to Eq. (67). The local velocity can increase (acceleration) or decrease (deceleration), but overall the kinetic energy $\mathcal{K}$ decreases with time in a closed system. Once dissipation takes place, i.e., $\mathrm{d}\mathbf{u}/\mathrm{d}t\neq 0$, the density at the interface cannot be assumed to be a constant; in this case, the density evolves with time. The density evolution is especially important when the fluid velocity is close to the speed of sound, i.e., high Mach number fluid. Note that in reality, dissipation cannot be avoided, since entropy always increases with time.
In Newton’s law, the time evolution of the velocity is defined as the “acceleration”, where the density evolution within the interface has been overlooked. Our concept provides a new understanding on the “acceleration”. The associated time evolution of the velocity corresponds to certain types of dissipation, coupled with the density evolution, which leads to the momentum balance. More generally speaking, we have divided the kinetic energy evolution into two parts:
$\frac{\mathrm{d}\mathcal{K}}{\mathrm{d}t}=\int_{\Omega}\bigg[\frac{\mathrm{d}(\rho\mathbf{u})}{\mathrm{d}t}\cdot\mathbf{u}+\rho\mathbf{u}\cdot\frac{\mathrm{d}\mathbf{u}}{\mathrm{d}t}\bigg]\mathrm{d}\Omega.$ (84)
The first part corresponds to the momentum balance. The second part is associated with the velocity dissipation. Overall, $\mathrm{d}\mathcal{K}/\mathrm{d}t\leq 0$, which is the key basis of the present finding.
VII Case Studies and Applications: Validation and Broader Implications
In this section, we apply the present theory to three representative cases: (A) sound propagation in air, (B) equation of state (EOS) for ideal gas, and (C) the motion of a water-air interface in a tube, to demonstrate the validity of the proposed theory. In case A, we derive a general formulation for the sound speed that is consistent with previous perturbation analysis. In case B, we show that the present density evolution equation in steady state generalizes the EOS of ideal gas. In case C, we demonstrate the validity of Bernoulli’s law, as well as its generalization accounting for the density jump at the water-air interface.
VII.1 Sound speed
When a sound wave propagates through air, the local density of the air varies with time—a hallmark of compressible flow. In an ideal gas, sound transmission occurs through molecular collisions. Over the timescale of these collisions, entropy remains constant (isentropic condition), implying no dissipation in the sound speed $c$ (i.e., $\mathrm{d}c/\mathrm{d}t=0$). Consequently, momentum evolution arises solely from density variations at the wavefront:
$c\frac{\mathrm{d}\rho}{\mathrm{d}t}=-\frac{\mathrm{d}p}{\mathrm{d}x}-\eta\frac{\mathrm{d}^{2}c}{\mathrm{d}x^{2}}.$ (85)
We postulate that sound propagation in a gas is driven by local variations in air density, not by changes in fluid velocity. Under the condition $c=\mathrm{d}x/\mathrm{d}t$, we derive the following equation at the wavefront when assuming isentropic condition ($\eta\approx 0$) via omitting the viscous term:
$c^{2}=-\frac{\mathrm{d}p}{\mathrm{d}\rho}.$ (86)
Using the pressure definition, $p=f-\sum_{i}\mu_{i}\phi_{i}$, we derive a general expression for the sound speed:
$c^{2}=\frac{\mathrm{d}(f-\sum_{i}\mu_{i}\phi_{i})}{\mathrm{d}\rho},\ \text{or},\ c=\sqrt{\frac{\mathrm{d}(f-\sum_{i}\mu_{i}\phi_{i})}{\mathrm{d}\rho}}=\sqrt{\frac{\mathrm{d}\Psi}{\mathrm{d}\rho}},$ (87)
which is consistent with the results of the perturbation analysis Landau and Lifshitz (2013). Note that two conditions have been applied to derive the sound speed: $\delta q\approx 0$ (see appendix) and $\eta\approx 0$, known as isentropic condition (adiabatic and reversible). For one component fluid, the Landau potential is expressed as $f-\sum_{i}\mu_{i}\phi_{i}=f$, where the free energy $f$ is written as $f=\theta-Ts$. In the case of isentropic condition, the free energy is mainly contributed by the internal energy, which consists of microscopic kinetic energy $\varrho$ and microscopic potential energy $\xi$,
$\theta=\varrho+\xi.$ (88)
The microscopic kinetic energy arising from the thermal motion of the gas molecules is expressed as $\varrho=\rho\bar{v}^{2}$, where $\bar{v}$ is the mean thermal velocity of the gas molecules. Performing its derivative to the density according to Eq. (87), we obtain the sound speed as
$c=\sqrt{\bar{v}^{2}+\frac{\mathrm{d}\xi}{\mathrm{d}\rho}}.$ (89)
For ideal gas, there is no interaction between gas molecules, and there is only the microscopic kinetic energy, namely, $\xi=0$, resulting in
$c\approx\bar{v}.$ (90)
This leads to an important conclusion that the sound speed in an ideal gas can be approximated by the mean thermal motion velocity of the gas molecules. For example, in air, the sound speed at room temperature and 1 bar is about 340 m/s, which is quite close to mean thermal velocity of gas molecules estimated via $\sqrt{R_{g}T/M}\approx 300$ m/s ($R_{g}$: universal gas constant, $M$: molecular weight). It must be noted that the above calculation only considers the kinetic energy and overlooks the interaction energy between gas molecules. The latter one cannot be neglected in real gas, solid, and liquid.
We further demonstrate that the density evolution equation is also consistent with the dispersion relation of sound propagation. When a sound wave propagates through air, the local air density varies temporally, which may be expressed as:
$\rho(x,t)=\rho_{0}\exp[-i(kx-\omega t)],$ (91)
where $k$ represents the wave number, $\omega$ depicts the frequency, and $\rho_{0}$ is a constant. Substituting the density wave into Eq. (85), the density evolution equation replicates the dispersion relation of sound wave
$\lambda=2\pi c/\omega.$ (92)
Here, $\lambda=2\pi/k$ is the wavelength of sound speed.
VII.2 Equation of the state (EOS)
In the steady state, the density evolution for inviscid fluids reduces to
$\rho u^{2}\frac{1}{\rho}\frac{\mathrm{d}\rho}{\mathrm{d}x}=-\frac{\mathrm{d}p}{\mathrm{d}x},$ (93)
where we have added a constant of unity: $\frac{\rho}{\rho}=1$, to the left hand side. When comprehending $u$ as the microscopic velocity, the kinetic energy $\rho u^{2}$ should equal to the thermal motion energy, $R_{g}T/v_{m}$ for ideal gas ($v_{m}$: molar volume). Integrating from a reference state with density $\rho_{0}$ and pressure $p_{0}$ to the state to be solved, Eq. (93) is rewritten as
$\frac{R_{g}T}{v_{m}}\int_{\rho_{0}}^{\rho}\frac{1}{\rho}\mathrm{d}\rho=-\int_{p_{0}}^{p}\mathrm{d}p.$ (94)
Replacing $v_{m}$ by $V/n$ ($V$: volume of the gas, $n$: number of mole), we obtain the following EOS
$nR_{g}T(\ln\rho-\ln\rho_{0})=(p_{0}-p)V.$ (95)
When choosing $p_{0}=0$ when $\ln\rho_{0}=0$, we obtain
$nR_{g}T\ln\rho=-pV=(f-\sum_{i}\mu_{i}\phi_{i})V=\Psi V,$ (96)
which is the EOS of ideal gas with a factor, $\ln\rho$. Note that the thermodynamic pressure $\Psi$ should be used instead of $p$. The term $\ln\rho$ is analogous to the entropy contribution, and this result is consistent with previous work on the free energy formulation of a water-gas system Ledesma-Aguilar et al. (2014). By dividing both sides of Eq. (96) by the molar volume, we obtain $\Psi\sim\rho\ln\rho$, as used in Ref. Ledesma-Aguilar et al. (2014).
VII.3 Water-air interface
As a further application, we consider a water-air system to demonstrate the present theory. Following the van der Waals-Cahn diffuse interface approach, the free energy functional for a binary system takes the form:
$\mathcal{F}=\int_{\Omega}\bigg[\frac{\gamma}{\epsilon}\phi^{2}(1-\phi)^{2}+\gamma\epsilon(\nabla\phi)^{2}\bigg]\mathrm{d}\Omega.$ (97)
The first term is a double-well potential with two local minima corresponding to the bulk water and bulk air phases. Interpreting $\phi$ as the volume concentration of water, $\phi=1$ represents pure water phase and $\phi=0$ corresponds to the pure gas phase. The second term accounts for the gradient energy density, originating from a Taylor expansion of the energy associated with concentration gradients at the interface Cai et al. (2024). The parameters $\gamma$ and $\epsilon$ control the interfacial tension and width, respectively. For simplicity, we set $\gamma=1$ and $\epsilon=1$ without loss of generality.
Figure 5: Analytical solution of the equation system, Eq. (98)-Eq. (102) in 2D for a moving flat water-air interface in a tube: (a) Volume concentration $\phi(x)$. (b) Pressure $p(x)$ and fluid velocity $u(x)$, with $u_{0}=0.025$ and $a=0.33$. (c) Density $\rho(x)/\rho_{w}$, where $\rho_{w}$ is the density of bulk water. Square symbols represent DFT results from Refs. Pezzotti et al. (2018); Creazzo and Luber (2024); the dashed line shows a linear interpolation of the density. (d) Sum of the pressure energy and kinetic energy, $p+\rho u^{2}$.
To minimize the free energy plus pressure and kinetic energies, the time evolution equations of the system read
$\frac{\partial\phi}{\partial t}+\mathbf{u}\cdot\nabla\phi=\nabla\cdot\varsigma\nabla\mu,$ (98)
$\mu=2\phi(1-\phi)-2\nabla^{2}\phi,$ (99)
$\mathbf{u}\frac{\mathrm{d}\rho}{\mathrm{d}t}=\mathbf{u}\bigg(\frac{\partial\rho}{\partial t}+\mathbf{u}\cdot\nabla\rho\bigg)=-\nabla p,$ (100)
$\frac{\mathrm{d}\mathbf{u}}{\mathrm{d}t}=\frac{\partial\mathbf{u}}{\partial t}+\mathbf{u}\cdot\nabla\mathbf{u}=-\tau_{0}\mathbf{u},$ (101)
$p=-\phi^{2}(1-\phi)^{2}-(\nabla\phi)^{2}+\mu\phi.$ (102)
The system comprises five equations for five unknowns: $(\phi,\mu,\rho,\mathbf{u},p)$. With appropriate boundary conditions, it admits a unique solution. In the following, we derive an analytical solution in 2D for a flat water-air interface moving towards the gas inside a tube with full slip condition (i.e., no wall friction, corresponding to a Young’s contact angle of 90o).
The analytical solution reads
$\phi(x)=\frac{1}{2}[1-\tanh(x)],$ (103)
$\mu(x)=0,$ (104)
$u_{0}^{2}\frac{\mathrm{d}\rho}{\mathrm{d}x}=-\frac{\mathrm{d}p}{\mathrm{d}x},$ (105)
$u(x)=u_{0}\exp(-ax),$ (106)
$p(x)=-\phi^{2}(1-\phi)^{2}-\frac{1}{4}[1-\tanh^{2}(x)]^{2}.$ (107)
The solution fulfils the boundary conditions
  • (B1)
    $\phi=1$ when $x\rightarrow$-$\infty$; $\phi=0$ when $x\rightarrow+\infty$.
  • (B2)
    $\frac{\mathrm{d}\phi}{\mathrm{d}x}=0$ when $x\rightarrow$-$\infty$ and $x\rightarrow+\infty$, i.e., no gradient of composition in the bulk water and gas.
  • (B3)
    According to (B1) and (B2), the boundary condition of the chemical potential is 0, i.e., $\mu(x)=0$ everywhere.
  • (B4)
    The fluid velocity in the water far from the interface is a constant, $u_{w}=u_{0}$. In the gas far from the interface, the fluid velocity is zero, $u_{a}=0$.
  • (B5)
    The bulk pressure of the water ($p_{w}$) and the bulk pressure of the gas ($p_{a}$) both are the same, fulling the Young-Laplace relation for a flat water-air interface.
The above solution is illustrated in Fig. 5 and indicates the following results:
  • (i)
    In the bulk water and bulk gas phases, we have $p_{w}+\frac{1}{2}\rho_{w}u_{w}^{2}=C_{1}$ and $p_{a}+\frac{1}{2}\rho_{g}u_{a}^{2}=C_{2}$, where $C_{1}$ and $C_{2}$ are two distinct constants. These results validate Bernoulli’s law. The difference in these two constants is proportional to the density difference $(\rho_{w}-\rho_{a})$. This energy jump has not been captured by Bernoulli’s law.
  • (ii)
    The pressure across the water-air interface first decreases from a constant value $p_{w}=0$ until the position $\phi(x=0)=0.5$ and then increases to a constant value $p_{a}=0$ in the gas phase. The space distribution is symmetric for a flat interface. For a curved water-air interface, Eq. (98) reproduces the Young-Laplace equation,
    $\Psi_{w}-\Psi_{a}=\sigma_{wa}\varGamma,$ (108)
    which has been already discussed in Ref. Wang et al. (2023) ($\sigma_{wa}$: surface tension; $\varGamma$: mean curvature).
  • (iii)
    The fluid velocity decays from a constant value in the bulk water to zero far from the interface in the gas. The spatial decay is linear when $\tau_{0}$ is a positive constant, as demonstrated by the steady state equation for fluid velocity:
    $u\frac{\mathrm{d}u}{\mathrm{d}x}=-\tau_{0}u.$ (109)
    As given by Eq. (106), the decay is exponential when $\tau_{0}$ is proportional to $u$, namely, $\tau_{0}=au$, where $a$ is a positive constant. The decay can also be Newtonian type
    $u\frac{\mathrm{d}u}{\mathrm{d}x}=\eta\frac{\mathrm{d}^{2}u}{\mathrm{d}x^{2}},$ (110)
    which will not be discussed in this work.
  • (iv)
    The density profile across the water-air interface is obtained via solving the ordinary differential equation, Eq. (105) with the boundary condition $\rho=\rho_{w}$ when $x\rightarrow$-$\infty$ and $\rho=\rho_{a}$ when $x\rightarrow+\infty$. Due to the special property of the pressure that the derivative $\mathrm{d}p/\mathrm{d}x$ is positive for $x\in(-\infty,0)$ and the derivative $\mathrm{d}p/\mathrm{d}x$ is negative for $x\in(0,\infty)$, the density first increases from $\rho=\rho_{w}$ to a maximum value at $x=0$ and then decreases to $\rho=\rho_{a}$ in the gas phase. The occurrence of a density peak within the water-air interface is fully consistent with recent results of the density functional theory (DFT) Pezzotti et al. (2018); Creazzo et al. (2020); Creazzo and Luber (2024). A change in the formulation of the velocity decay cannot remove the peak, but can affect the peak value as well as the peak width. It must be noted the conventional interpolation approach, $\rho=\rho_{w}\phi+\rho_{a}(1-\phi)$ can only lead to a monotonic decrease of the density from $\rho_{w}$ to $\rho_{a}$, contradicting with the DFT results.
VIII Conclusion
In summary, we have proposed an alternative theory, distinct from the Navier-Stokes and Euler equations, to address the high-density-ratio problem in fluid dynamics. Our approach is grounded in the fundamental principle of energy minimization. A noteworthy outcome of this work is the derivation of a novel density evolution equation
$\mathbf{u}\dot{\rho}=-\nabla\cdot[(p\mathbf{I}+\underline{\underline{\bm{\Theta}}})]-\rho\nabla\cdot[\tau_{1}(\nabla\bm{\zeta}+\nabla\bm{\zeta}^{T})+\tau_{2}(\nabla\cdot\bm{\zeta})\mathbf{I}]$ (111)
for the high density ratio system, such as water-air. From this equation, we see that the density evolution or the local excess volume is affected by two factors:
  • the local pressure $p$ and stress tensor $\underline{\underline{\bm{\Theta}}}$;
  • the viscous dissipation.
The latter one provides a supplement to the Navier-Stokes equation that the density can be altered by the viscous dissipation. The validity of our density evolution equation has been demonstrated for inviscid fluids. Specially, our derivation leads to a generalization of Bernoulli’s law. In addition, we have shown that the current density evolution is consistent with
  • sound speed in fluids;
  • dispersion relation of sound wave;
  • EOS of ideal gas;
  • DFT results for non-monotonic density across the water-air interface.
We expect that the present model can be used for immiscible fluids with arbitrary density ratios.
Acknowledgements.
This research is supported by VirtMat project of the Helmholtz association, as part of the programme “MSE-materials science and engineering” number-43.31.01. F.W. proposed the concept, carried out the theoretical derivations, and prepared the manuscript. B.N. contributed to securing the funding.
Appendix A Gibbs-Duhem equation and the kinetic energy
In the case of incompressible fluids, a single variable—either $\phi$ or $\rho$—is sufficient to characterize the state of the system. For compressible fluids with nearly identical densities, a single variable likewise suffices to adequately describe the system, as demonstrated in Section III. Here, we illustrate the Gibbs–Duhem equation by taking $\phi$ as the state variable. According to the first law of thermodynamics, the change in the internal energy density $\theta$ of a unit volume element $v$ can be expressed as
$\mathrm{d}\theta=\delta q-p\mathrm{d}v+\sum_{i}\mu_{i}\mathrm{d}\phi_{i}.$ (112)
The first term corresponds to heat transfer, while the second and third terms represent mechanical and chemical work, respectively. Typically, the timescale associated with heat transfer, represented by $\delta q$, is much longer than that of mass transport, as thermal diffusivity exceeds mass diffusivity by at least three orders of magnitude. Consequently, we assume an isothermal condition on the timescale of mass transport and adopt $\delta q\approx 0$ in the current work. Nonetheless, the temperature evolution equation is derived in Section V to provide a complete description of the system.
Combining Eq. (112) with the total derivative $\mathrm{d}\theta=-p\mathrm{d}v-v\mathrm{d}p+\sum_{i}\mu_{i}\mathrm{d}\phi_{i}+\sum_{i}\phi_{i}\mathrm{d}\mu_{i}$, we arrive at the Gibbs–Duhem equation:
$v\mathrm{d}p+\sum_{i}\phi_{i}\mathrm{d}\mu_{i}=0,$ (113)
which has been often used to the formulate the pressure force for a unit volume
$\nabla p=\sum_{i}\phi_{i}\nabla\mu_{i}$ (114)
for the momentum evolution.
We note that a limitation of the above derivation is the omission of the kinetic energy, $\rho\mathbf{u}^{2}$ due to convection. The classical Gibbs–Duhem relation assumes thermodynamic equilibrium at each time step—that is, a quasi-static process—and therefore does not account for the momentum balance associated with kinetic energy. Another limitation arises in systems with large density ratios, where two variables, $\rho$ and $\phi$, are required to fully describe the change in system energy. In this context, the quintessential Gibbs–Duhem relation considers only the mechanical/chemical work associated with $\mathrm{d}\phi$, while the contribution from $\mathrm{d}\rho$ is neglected. A more complete treatment should consider the total energy, including internal energy ($\theta$), kinetic energy ($\rho\mathbf{u}\cdot\mathbf{u}$), and thermodynamic pressure ($p$), i.e., $\mathrm{d}(\theta+p+\rho\mathbf{u}\cdot\mathbf{u})$, rather than only $\mathrm{d}\theta$ as in the quasi-static approximation. When the contribution of the mixing entropy $s$ is considered, the internal energy $\theta$ should be replaced by the free energy density $f=\theta-Ts$. One illustrative example is that minimizing the sum of the thermodynamic pressure and kinetic energy naturally leads to momentum conservation:
$\mathrm{d}(p+\rho\mathbf{u}\cdot\mathbf{u})=\mathrm{d}p+\mathrm{d}(\rho\mathbf{u})\cdot\mathbf{u}+\rho\mathbf{u}\cdot\mathrm{d}\mathbf{u}.$ (115)
In this expression, the first two terms contribute to momentum conservation: $\mathrm{d}(\rho\mathbf{u})/\mathrm{d}t=-\nabla p$ by reformulating $\mathrm{d}p$ as $\mathrm{d}p/\mathrm{d}t=(\mathrm{d}p/\mathrm{d}\mathbf{r})\cdot(\mathrm{d}\mathbf{r}/\mathrm{d}t)=\nabla p\cdot\mathbf{u}$). The last term in Eq. (115) represents the dissipation of velocity. The comprehensive energy formulation encompassing the free energy, pressure energy, kinetic energy, and wall free energy is expressed in Eq. (52). The detailed derivation of the corresponding energy minimization, which represents a key result of the present work, is presented in Section VI.

### v3

@@
 In this work, we explore the high-density ratio problem through the physical principle of energy minimization.
-We propose a generalized theoretical framework for the evolution equations of fluid dynamics and derive a novel kinetic equation governing density dynamics at the water-air interface.
-The resulting density evolution equation yields three key contributions: (A): A universal formulation for sound speed; (B): A generalized equation of state for an ideal gas; (C): A generalization of Bernoulli’s law, extending the conservation of potential and kinetic energy.
+We propose a generalized theoretical framework for the evolution equations of fluid dynamics and derive a novel kinetic equation governing density dynamics at the fluid-fluid interface.
+The resulting density evolution equation yields three key contributions: (A): A universal formulation for sound speed; (B): A generalized equation of state for fluids; (C): A generalization of Bernoulli’s law, extending the conservation of potential and kinetic energy.
 For each contribution, a case study and application is presented in Section VII.
 The rest of the paper is structured as follows.
-In Section II, we provide a brief overview of two high-density ratio models from the literature, clarify the distinction between the total time derivative and the mass conservative derivative, and examine the validity of the conventional continuity equation.
+In Section II, we provide a brief overview of two high-density ratio models from literature, clarify the distinction between the total time derivative and the mass conservative derivative, and examine the validity of the conventional continuity equation.
 Section III outlines the setup of our study and introduces a new concept of volume concentration.
@@
 Section V applies this principle to derive the classical fluid dynamics equations for uniform systems.
-Building on these results, Section VI extends the derivation to non-uniform systems, derives the energy dissipation, and formulates a novel kinetic equation for density.
-In Section VII, we demonstrate the application of our theory by deriving the sound speed, the equation of state (EOS) for an ideal gas, and the density profile across a water-air interface.
+Building on these results, Section VI extends the derivation to non-uniform systems, derives the free energy dissipation, and formulates a novel kinetic equation for density.
+In Section VII, we demonstrate the application of our theory by deriving the sound speed, the equation of state (EOS) for fluids, and the density profile across a water-air interface.
 Finally, Section VIII concludes the paper with a summary of our findings.
@@
 (1997); Fedkiw et al.
-(1999) (which will be discussed in Section II C) or on the linear interpolation of the density Fakhari et al.
+(1999) (which will be discussed in Section II C and Section VI D) or on the linear interpolation of the density Fakhari et al.
 (2017); Shukla et al.
@@
 This redistribution of excess volume implies that the local partial densities of water and gas must vary accordingly, since the individual masses of the two components within the volume element $V$ remain constant.
-We define the volume composition as
+We define the volume composition including the information of excess volume as
 $\phi_{w}=\frac{v_{w}^{\prime}}{V}\quad\text{and}\quad\phi_{a}=\frac{v_{a}^{\prime}}{V}.$ (29)
@@
 $\rho=\rho_{w}^{\prime}\frac{v_{w}^{\prime}}{V}+\rho_{a}^{\prime}\frac{v_{a}^{\prime}}{V}=\rho_{w}^{\prime}\phi+\rho_{a}^{\prime}(1-\phi).$ (30)
+Note that both $\rho$ and $\phi$ encode the information associated with the excess volume.
 The partial densities $\rho_{w}^{\prime}$ and $\rho_{a}^{\prime}$ are implicitly influenced by the excess volume and are therefore no longer constant, as they vary with the redistribution of the excess volume.
@@
 (2004); Lowengrub and Truskinovsky (1998), in which a single phase order parameter $\phi$ is used to characterize the system’s state and evolution.
-This well-established approach is commonly referred to as the phase-field model or Cahn–Hilliard model.
+This well-established approach Ridl and Wagner (2018) is commonly referred to as the phase-field model or Cahn–Hilliard model.
 Another classical example is Euler’s formulation for density evolution, where the standard continuity equation, Eq.
@@
 $\frac{\mathrm{d}u}{\mathrm{d}t}=\nabla\cdot\tau_{0}\nabla\Phi=\nabla\cdot\tau_{0}\nabla\frac{\mathrm{d}e}{\mathrm{d}u},$ (34)
-where $\tau_{0}$ is a positive constant. Substituting Eq. (34) into Eq. (33), integrating by parts, and using the no-flux boundary condition, we obtain
+which is consistent with standard theory of diffusion equation.
+The coefficient $\tau_{0}$ is a positive constant.
+Substituting Eq.
+(34) into Eq.
+(33), integrating by parts, and using the no-flux boundary condition, we obtain
 $\frac{\mathrm{d}E}{\mathrm{d}t}=\int_{\Omega}\frac{\mathrm{d}e}{\mathrm{d}u}\frac{\mathrm{d}u}{\mathrm{d}t}\mathrm{d}\Omega=\int_{\Omega}\frac{\mathrm{d}e}{\mathrm{d}u}\nabla\cdot\tau_{0}\nabla\Phi\mathrm{d}\Omega=-\int_{\Omega}\tau_{0}(\nabla\Phi)^{2}\mathrm{d}\Omega\leq 0.$ (35)
@@
 Thus, the coefficient $\tau_{0}$ establishes a connection between the purely dynamic parameter $\mathcal{D}$ (independent of conservative energy) and the energy-dependent term $\frac{\mathrm{d}^{2}e}{\mathrm{d}u^{2}}$ via the relation $\tau_{0}=\mathcal{D}/\frac{\mathrm{d}^{2}e}{\mathrm{d}u^{2}}$.
-In other words, $\tau_{0}$ represents the ratio of non-conservative to conservative energy contributions to the kinetics of the system.
 The dissipation equation for non-conserved variables can be written in the steepest descendent approach:
@@
 $\mathcal{K}$ $=\int_{\Omega}k\mathrm{d}\Omega=\int_{\Omega}\rho\mathbf{u}\cdot\mathbf{u}\mathrm{d}\Omega.$ (41)
-The free energy density depends on the volume concentration $\phi$ and temperature $T$, and may be expressed as $f=\theta-Ts$, where $\theta$ denotes the internal energy and $s\sim\phi\ln\phi$ represents the mixing entropy, which is distinct from the thermal (heat) entropy.
+The free energy density depends on the volume concentration $\phi$ and temperature $T$, and may be expressed as $f=\theta-Ts$, where $\theta$ denotes the internal energy and $s$ represents the entropy.
 The total time derivative of the total energy reads
@@
 $\frac{\mathrm{d}\phi}{\mathrm{d}t}=\nabla\cdot M\nabla\frac{\partial f}{\partial\phi}=\nabla\cdot M(\partial^{2}f/\partial\phi^{2})\nabla\phi=\nabla\cdot D\nabla\phi,$ (46)
-where the mass diffusivity $D$ is defined in terms of the mobility $M$ as $D=M(\mathrm{d}^{2}f/\mathrm{d}\phi^{2})$.
+where the mass diffusivity $D$ is defined in terms of the mobility $M$ as $D=M(\partial^{2}f/\partial\phi^{2})$.
 When considering that there is no velocity dissipation, e.g., inviscid fluids, we obtain the momentum balance equation according to the second term in Eq. (44)
@@
 $\mathcal{E}=\mathcal{F}+\mathcal{P}+\mathcal{K}+\mathcal{W}.$ (52)
-In contrast to the uniform system, the free energy of a non-uniform system depends not only on $\phi$ but also on $\nabla\phi$, as derived by Cahn Cahn and Hilliard (1958) and van der Waals Rowlinson (1979), known as van der Waals-Cahn theory.
+This is a standard approach in the literature, except for the inclusion of the pressure term, $\mathcal{P}$. The motivation for introducing $\mathcal{P}$ will be presented below.
 The integration over the local free energy density $e(\bm{\phi},\nabla\bm{\phi})$ leads to
-$\mathcal{F}(\bm{\phi},\nabla\bm{\phi})=\int_{\Omega}e(\bm{\phi},\nabla\bm{\phi})d\Omega,$ (53)
-$e(\bm{\phi},\nabla\bm{\phi})=f(\bm{\phi})+\sum_{i=1}^{K}\frac{1}{2}\kappa_{i}(\nabla\phi_{i})^{2},$ (54)
-where $f(\bm{\phi})$ depicts the bulk free energy density and $\kappa_{i}$ represents the gradient energy coefficient which is related to the interfacial tension of the $\alpha$-$\beta$ interface, $\sigma_{\alpha\beta}$.
-Figure 3: Origin of the pressure energy for uniform and nonuniform systems.
-For a uniform system (a), the pressure at any position is the same, which is taken as a reference value.
-In a non-uniform system with density variation (b), the pressure is generally non-uniform due to the asymmetric collisions of different sized molecules.
-Blue and red spheres depict two molecules or atoms with distinct sizes, leading to a variation of the density in space.
-The arrows sketch the thermal motion of the molecules/atoms.
-The $\mathcal{P}=\int_{\Omega}pd\Omega$ term, often overlooked previously and introduced here, describes the pressure energy. We formulate the pressure energy density as Chen and Doolen (1998)
-$-p=e-\sum_{i=1}^{K}\mu_{i}\phi_{i},$ (55)
-where $\mu_{i}$ is the chemical potential of species $i$, the formulation of which will be derived in the following via the energy dissipation principle.
-Herein, an essential consideration is that the pressure $p$ is always associated with the free energy $\mathcal{F}$.
-A physical interpretation of the pressure is shown in Fig.
-3.
-A homogeneous mixture of two distinct species (Fig.
-3a) leads to a uniform pressure $p_{0}(\bm{r})=\text{const},\ \forall\bm{r}\in\Omega$, which can be taken as a reference value.
-However, when the two species are not homogeneously mixed (Fig.
-3b), the intrinsic pressure is non-uniform in space; the non-uniform pressure is caused, for example, by the collisions of different molecules with distinct sizes due to the asymmetry in space.
-In the special case of a homogeneous mixture, the pressure defined in Eq.
-(55) is consistent with the Landau potential (sometimes called the thermodynamic pressure Wang et al.
-(2023)), given by
-$\Psi=f(\bm{\phi})-\sum_{i=1}^{K}(\partial f/\partial\phi_{i})\phi_{i}.$ (56)
+$\mathcal{F}(\bm{\phi},\nabla\bm{\phi})=\int_{\Omega}e(\bm{\phi},\nabla\bm{\phi})\mathrm{d}\Omega.$ (53)
+By decomposing the entropy $s$ into thermal entropy $s_{T}$, component-mixing entropy $s_{\phi}$, and velocity entropy $s_{u}$, we write the free energy density as
+$e=\theta-Ts=\theta-T(s_{T}+s_{\phi}+s_{u})=e_{0}-Ts_{u},$ (54)
+where $\theta$ represents the internal energy. The physical meaning of the entropies $s_{\phi}$ and $s_{u}$ is as follows:
+  • The component mixing entropy $s_{\phi}$ accounts for the mixing excess volume in the diffusion equation, thereby leading to compressibility.
+  • The velocity entropy $s_{u}$ is associated with viscous effects, resulting in velocity dissipation.
+The free energy $e_{0}=f-T(s_{T}+s_{\phi})$ does not contain the contribution of the velocity entropy.
+Here, $f=f_{0}+\sum_{i=1}^{K}\kappa_{i}(\nabla\phi_{i})^{2}$ depicts the van der Waals-Cahn free energy Cahn and Hilliard (1958); Rowlinson (1979) and $\kappa_{i}$ represents the gradient energy coefficient.
+It should be noted that the entropy increase of $s_{\phi}$ and $s_{u}$ does not necessarily result in heat release or a temperature change of the same magnitude as that associated with thermal entropy.
+For instance, mass diffusion can lead to an increase in $s_{\phi}$ in an approximately isothermal manner, since the mass diffusivity is much smaller than the thermal diffusivity.
+Similarly, dissipation of fluid velocity can increase $s_{u}$ nearly isothermally, as the kinematic viscosity of water is one magnitude smaller than its thermal diffusivity.
+For these physical reasons, we assume isothermal conditions over the time scale of water–mass diffusion in the present work.
 The $\mathcal{K}$ term denotes the kinetic energy of the fluid, which is formulated as
-$\mathcal{K}=\int_{\Omega}\bm{\zeta}\cdot\mathbf{u}d\Omega,$ (57)
+$\mathcal{K}=\int_{\Omega}\bm{\zeta}\cdot\mathbf{u}\mathrm{d}\Omega,$ (55)
 where $\bm{\zeta}$ is the linear momentum density expressed as $\bm{\zeta}=\rho\mathbf{u}$.
+The $\mathcal{P}=\int_{\Omega}p\mathrm{d}\Omega$ term, often overlooked previously and introduced here, describes the pressure energy.
+In existing works, the space gradient of pressure is usually employed as the driving force for the fluid motion in the Navier-Stokes equation by using the isothermal Gibbs-Duhem relation
+$\mathrm{d}p+\sum_{i=1}^{K}\phi_{i}\mathrm{d}\mu_{i}=0,$ (56)
+where $\mu_{i}$ is the chemical potential of species $i$, defined as $\mu_{i}=\delta e/\delta\phi_{i}$.
+Here, $\delta$ denotes the functional derivative.
+However, we note that there are two major limitations when using the gradient of pressure $\nabla p$ based on $\mu_{i}=\delta e/\delta\phi_{i}$ and Eq.
+(56) to derive the momentum evolution.
+  • First, the classical Gibbs–Duhem relation, Eq.
+(56), is derived under a quasi-static assumption, in which the velocity-dependent entropy $s_{u}$ and the associated kinetic energy are not taken into account.
+Consequently, when $\nabla p$ is used as given in Eq.
+(56), the role of viscosity in free-energy dissipation or in maximizing the velocity entropy is not clearly defined.
+One cannot simply introduce a viscous term into the momentum evolution in an ad hoc manner without accounting for its fundamental connection to entropy.
+  • Second, the space derivative $\nabla$ and the functional derivative $\delta$ (or $\mathrm{d}$) are not mathematically compatible in most existing works except few references Ridl and Wagner (2018) (see Appendix B).
+To solve these two limitations, we derive a modified Gibbs-Duhem relation as follows.
+The total derivative of the kinetic energy reads, $\delta k=\mathbf{u}\cdot\delta(\rho\mathbf{u})+\rho\mathbf{u}\cdot\delta\mathbf{u}$.
+The first part $\mathbf{u}\cdot\delta(\rho\mathbf{u})$ is the inertial work which can be absorbed into the first law of thermodynamics in addition to the chemical work $\mu_{i}\delta\phi_{i}$ and mechanical work $p\delta v$.
+As will be shown in Section VIB, this term, $\mathbf{u}\cdot\delta(\rho\mathbf{u})$ combing with pressure results in the momentum evolution.
+The second part $\rho\mathbf{u}\cdot\delta\mathbf{u}$ is left into the modified isothermal Gibbs-Duhem relation
+$\delta p+\sum_{i=1}^{K}\phi_{i}\delta\mu_{i}+\rho\mathbf{u}\cdot\delta\mathbf{u}=0,$ (57)
+as the spirit of Gibbs-Duhem relation is that the total derivative of the energy minus the work associated with the first law of thermodynamics is zero.
+As $\rho\mathbf{u}\cdot\delta\mathbf{u}=-T\delta s_{u}$ (see Section VIB), we integrate Eq.
+(57) and obtain a general definition of the pressure as (see Appendix A)
+$p=e-\sum_{i=1}^{K}\mu_{i}\phi_{i}+Ts_{u}=e_{0}-\sum_{i=1}^{K}\mu_{i}\phi_{i}.$ (58)
+This result provides a clear understanding of the pressure and is fully consistent with previous works Chen and Doolen (1998).
+It must be noted that the pressure $p$ in Eq.
+(58) does not include the velocity-dependent entropy, as $e+Ts_{u}$ results in a $s_{u}$ independent energy $e_{0}$.
 The $\mathcal{W}$ term in Eq.
-(52) represents the wall free energy, which is integrated over the wall free energy density $f_{w}$ on the boundary $\Gamma$ as $\mathcal{W}=\int_{\Gamma}f_{w}d\Gamma$.
-The formulation of the wall free energy density $f_{w}$ has been comprehensively discussed elsewhere Wang and Nestler (2024); Wang et al.
-(2023); Wang and Nestler (2021) and is not the scope of this work.
+(52) represents the wall free energy, which is integrated over the wall free energy density $f_{w}$ on the boundary $\Gamma$ as $\mathcal{W}=\int_{\Gamma}f_{w}\mathrm{d}\Gamma$.
+The wall free energy density $f_{w}$ is responsible for the wetting phenomenon, which has been comprehensively discussed elsewhere Wang and Nestler (2024); Wang et al.
+(2023); Wang and Nestler (2021).
 In the following, we derive the equilibrium condition and the kinetic equations based on the principle of energy minimization, namely
@@
 By using the chain rule and noting that the functional $\mathcal{F}$ has two arguments, $\bm{\phi}$ and $\nabla\bm{\phi}$, the total time derivative of the potential free energy $\mathcal{F}$ reads
-$\dot{\mathcal{F}}(\bm{\phi},\nabla\bm{\phi})=\int_{\Omega}\sum_{i=1}^{K}\frac{\partial e}{\partial\phi_{i}}\frac{\mathrm{d}\phi_{i}}{\mathrm{d}t}+\frac{\partial e}{\partial\nabla\phi_{i}}\cdot\frac{\mathrm{d}\nabla\phi_{i}}{\mathrm{d}t}d\Omega.$ (58)
-By using the vector calculus $\nabla\cdot(\nabla\phi_{i}\otimes\nabla\phi_{i})=\nabla^{2}\phi_{i}\nabla\phi_{i}+(\nabla\phi_{i}\cdot\nabla)\nabla\phi_{i}$, the definition of materials derivative $\dot{\nabla\phi_{i}}=\partial_{t}\nabla\phi_{i}+\mathbf{u}\cdot\nabla\nabla\phi_{i}$, and integration by parts with no-flux boundary condition for a closed system (more details can be found in Wang et al.
+$\dot{\mathcal{F}}(\bm{\phi},\nabla\bm{\phi})=\int_{\Omega}\sum_{i=1}^{K}\frac{\partial e}{\partial\phi_{i}}\frac{\mathrm{d}\phi_{i}}{\mathrm{d}t}+\frac{\partial e}{\partial\nabla\phi_{i}}\cdot\frac{\mathrm{d}\nabla\phi_{i}}{\mathrm{d}t}\mathrm{d}\Omega.$ (59)
+By using the vector calculus $\nabla\cdot(\nabla\phi_{i}\otimes\nabla\phi_{i})=\nabla^{2}\phi_{i}\nabla\phi_{i}+(\nabla\phi_{i}\cdot\nabla)\nabla\phi_{i}$, the definition of materials derivative $\dot{\nabla\phi_{i}}=\partial_{t}\nabla\phi_{i}+\mathbf{u}\cdot\nabla\nabla\phi_{i}$, and integration by parts with no-flux boundary condition (more details can be found in Wang et al.
 (2023); Zhang et al.
 (2024)), we obtain the following equations
-$\dot{\mathcal{F}}=\int_{\Omega}\sum_{i=1}^{K}\mu_{i}\dot{\phi}_{i}+(\nabla\cdot\underline{\underline{\bm{\Theta}}})\cdot\mathbf{u}d\Omega;$ (59)
-$\mu_{i}=\frac{\partial e}{\partial\phi_{i}}-\nabla\cdot\frac{\partial e}{\partial\nabla\phi_{i}};$ (60)
+$\dot{\mathcal{F}}=\int_{\Omega}\sum_{i=1}^{K}\mu_{i}\dot{\phi}_{i}+(\nabla\cdot\underline{\underline{\bm{\Theta}}})\cdot\mathbf{u}\mathrm{d}\Omega;$ (60)
+$\mu_{i}=\frac{\delta e}{\delta\phi_{i}}=\frac{\partial e}{\partial\phi_{i}}-\nabla\cdot\frac{\partial e}{\partial\nabla\phi_{i}};$ (61)
 $\underline{\underline{\bm{\Theta}}}=\sum_{i}\frac{\partial e}{\partial\nabla\phi_{i}}\otimes\nabla\phi_{i}.$ (62)
@@
 (60) defines a stress tensor $\underline{\underline{\bm{\Theta}}}$, which is responsible for the transformation of the potential free energy into the macroscopic kinetic energy Jacqmin (1999).
-This stress tensor is consistent with the Korteweg stress by evaluating the derivative via considering Eq.
-(54), namely $\partial_{\nabla\phi_{i}}e=\kappa_{i}\nabla\phi_{i}$.
+This stress tensor is consistent with the Korteweg stress by evaluating the derivative, $\partial_{\nabla\phi_{i}}e=\kappa_{i}\nabla\phi_{i}$.
 Noteworthily, a uniform chemical potential throughout the system cannot well define the thermodynamic equilibrium; the pressure evolution should be considered as well.
@@
 (63) and Eq.
-(63), we have used the chain rule and the definition of the velocity, $\dot{\mathbf{r}}=\mathbf{u}$.
+(64), we have used the chain rule and the definition of the velocity, $\dot{\mathbf{r}}=\mathbf{u}$ without considering relativity effect Landau and Lifshitz (2013).
+Note that $\mathrm{d}p/\mathrm{d}t=\partial p/\partial t+\mathbf{u}\cdot\nabla p$.
+For fluids with velocity much less than sound speed, $\partial p/\partial t\approx 0$.
 Summarizing Eq.
@@
 (64) and collecting all the $\cdot\mathbf{u}$ terms, we obtain the total time derivative of the system energy $\mathcal{E}$ as
-$\dot{\mathcal{E}}=\int_{\Omega}\sum_{i=1}^{K}\underbrace{\mu_{i}\dot{\phi_{i}}}_{\text{I}}+\underbrace{[\nabla\cdot(\underline{\underline{\bm{\Theta}}}+p\mathbf{I})+\dot{\bm{\zeta}}]}_{\text{II}}\cdot\mathbf{u}+\underbrace{\bm{\zeta}\cdot\dot{\mathbf{u}}}_{\text{III}}d\Omega.$ (64)
-The physical interpretation for the three parts of the energy evolution in Eq. (64) is sketched in Fig. 4.
-Figure 4: The energy dissipation mechanisms in Eq.
-(64).
-I: The diffusion via the exchange of the molecules and vacancies leads to the dissipation of the free energy $\mathcal{F}$.
-II: The gradient of the free energy $\mathcal{F}$ and the associated pressure $\mathcal{P}$ result in an increase in the momentum $\bm{\zeta}=\rho\mathbf{u}$.
-III: The difference in the velocity, density, and viscosity in distinct subdomains (dashed rectangle) gives rise to the dissipation of the velocity $\mathbf{u}$ caused by the ‘spring’ drag effect; the spring symbol denote the interaction between the atoms/molecules in different subdomains.
+$\dot{\mathcal{E}}=\int_{\Omega}\sum_{i=1}^{K}\underbrace{\mu_{i}\dot{\phi_{i}}}_{\text{I}}+\underbrace{[\nabla\cdot(\underline{\underline{\bm{\Theta}}}+p\mathbf{I})+\dot{\bm{\zeta}}]}_{\text{II}}\cdot\mathbf{u}+\underbrace{\bm{\zeta}\cdot\dot{\mathbf{u}}}_{\text{III}}\mathrm{d}\Omega.$ (65)
+Note that, at isothermal condition, the total energy is conserved when defining $\Lambda=\mathcal{E}+Ts_{\phi}+Ts_{u}$, where the entropy energies of component and velocity have been added.
+Within this framework, we write the time derivative of the total energy as
+$\frac{\mathrm{d}\Lambda}{\mathrm{d}t}=\int_{\Omega}\sum_{i=1}^{K}\bigg(\mu_{i}\frac{\mathrm{d}\phi_{i}}{\mathrm{d}t}+T\frac{\mathrm{d}s_{\phi_{i}}}{\mathrm{d}t}\bigg)+[\nabla\cdot(\underline{\underline{\bm{\Theta}}}+p\mathbf{I})+\dot{\bm{\zeta}}]\cdot\mathbf{u}+\bigg(\bm{\zeta}\cdot\frac{\mathrm{d}\mathbf{u}}{\mathrm{d}t}+T\frac{\mathrm{d}s_{u}}{\mathrm{d}t}\bigg)\mathrm{d}\Omega=0.$ (66)
 VI.2 Kinetic equations
@@
 $\dot{\phi_{i}}$ $=\frac{\mathrm{d}\phi_{i}}{\mathrm{d}t}=\frac{\partial\phi_{i}}{\partial t}+\mathbf{u}\cdot\nabla\phi_{i}=\nabla\cdot\varsigma_{i}\nabla\mu_{i}.$ (67)
-The coefficient $\varsigma_{i}$ is the mobility.
+The coefficient $\varsigma_{i}$ is the mobility. The physical meaning is that the diffusion leads to an increase in the mixing entropy $s_{\phi}$.
 The second part in Eq. (65), II, reveals the momentum conservation
@@
 The conservative force at the right hand side is responsible for the exchange of the potential energy $\mathcal{F}+\mathcal{P}$ with the kinetic energy $\mathcal{K}$.
-This result is nothing but Newton’s law of momentum balance.
+This result is nothing but momentum balance.
 Here, a more clear physical meaning of the force is presented; the force term consists of the pressure $p$ and the stress tensor $\underline{\underline{\bm{\Theta}}}$.
 The stress tensor is due to the gradient term in the potential energy $\mathcal{F}$, while the pressure is contributed by the associated free energy.
-Note that a viscous dissipation term cannot be directly added to the momentum conservation equation, as there is no quantity corresponding to “viscous energy”.
-The only term left in the time evolution of system energy, III, is $\int_{\Omega}\bm{\zeta}\cdot\dot{\mathbf{u}}\mathrm{d}\Omega$.
+Note that a viscous dissipation term cannot be added to the momentum conservation equation, as the pressure term does not contain the information of the velocity entropy [see Eq.
+(58)].
+The only term left in the time evolution of system free energy, III, is $\int_{\Omega}\bm{\zeta}\cdot\dot{\mathbf{u}}\mathrm{d}\Omega$.
 By applying the same dissipation principle as done for the scalar field in Eq.
@@
 $\dot{\mathbf{u}}=\nabla\cdot\tau_{1}(\nabla\bm{\zeta}+\nabla\bm{\zeta}^{T})+\nabla\cdot\tau_{2}(\nabla\cdot\bm{\zeta})\mathbf{I}.$ (69)
-Unlike the dissipation rule for the scalar variable $\phi_{i}$ with the associated potential $\nabla\mu_{i}$ (see Eq.
-(65)), three terms $\nabla\bm{\zeta}$, $\nabla\bm{\zeta}^{T}$, and $(\nabla\cdot\bm{\zeta})\mathbf{I}$ have to be taken into account for the dissipation of a vector $\mathbf{u}$.
+Unlike the dissipation rule for the scalar variable $\phi_{i}$ with the associated potential $\nabla\mu_{i}$ [see Eq.
+(67)], three terms $\nabla\bm{\zeta}$, $\nabla\bm{\zeta}^{T}$, and $(\nabla\cdot\bm{\zeta})\mathbf{I}$ have to be taken into account for the dissipation of a vector $\mathbf{u}$.
+The physical meaning of the velocity dissipation is the increase of the velocity entropy $s_{u}$.
+Within the framework of entropy, the third term, III, in Eq.
+(65) can be interpreted as
+$\bm{\zeta}\cdot\frac{\mathrm{d}\mathbf{u}}{\mathrm{d}t}=-T\frac{\mathrm{d}s_{u}}{\mathrm{d}t}.$ (70)
+The left hand side is responsible for the minimization of the kinetic energy, which is equivalent to the right hand side for the maximization of the velocity entropy.
 The right hand side of Eq. (69) is equivalent to Newton viscous stress tensor if we define the mobilities $\tau_{1}$ and $\tau_{2}$ as
@@
 The dissipation equation, Eq. (69), is essentially a diffusion equation applied to the vector $\mathbf{u}$, as evident from the following comparison:
-$\nabla\cdot\tau_{1}\nabla\bm{\zeta}=\nabla\cdot\tau_{1}\nabla(\partial\mathcal{K}/\partial\mathbf{u});\ \bm{\zeta}=\partial\mathcal{K}/\partial\mathbf{u};$ (69)
-$\nabla\cdot\varsigma\nabla\mu=\nabla\cdot\varsigma\nabla(\partial\mathcal{F}/\partial\mathbf{\phi});\ \mu=\partial\mathcal{F}/\partial\phi.$ (70)
+$\nabla\cdot\tau_{1}\nabla\bm{\zeta}=\nabla\cdot\tau_{1}\nabla(\partial k/\partial\mathbf{u});\ \bm{\zeta}=\partial k/\partial\mathbf{u};$ (72)
+$\nabla\cdot\varsigma\nabla\mu=\nabla\cdot\varsigma\nabla(\partial e/\partial\mathbf{\phi});\ \mu=\partial e/\partial\phi.$ (73)
 The linear momentum, $\bm{\zeta}$, plays a role analogous to that of the chemical potential, $\mu$.
@@
 $\frac{\mathrm{d}\mathbf{u}}{\mathrm{d}t}=-\tau_{0}\bm{\zeta},$ (74)
-where $\tau_{0}$ is a positive mobility. The non-conserved form is known as the Allen-Cahn type dissipation following the gradient descent path, being consistent with the Langevin equation.
-VI.3 The problem of the conventional kinetic energy formulation
-Conventional formulation for the kinetic energy is $\tilde{\mathcal{K}}=\int_{\Omega}\frac{1}{2}\rho\mathbf{u}^{2}\mathrm{d}\Omega$. In this case, the time derivative of the kinetic energy reads
-$\frac{\mathrm{d}{\tilde{\mathcal{K}}}}{\mathrm{d}t}=\int_{\Omega}\bigg(\frac{1}{2}\frac{\mathrm{d}\rho}{\mathrm{d}t}\mathbf{u}^{2}+\rho\mathbf{u}\cdot\frac{\mathrm{d}\mathbf{u}}{\mathrm{d}t}\bigg)\mathrm{d}\Omega.$ (72)
+The non-conserved form is known as the Allen-Cahn type dissipation following the gradient descent path, being consistent with the Langevin equation.
+VI.3 Density evolution equation and the generalization of Bernoulli’s law
+Eq. (68) subtracting Eq. (69) with the relation $\dot{\bm{\zeta}}=\dot{\rho}\mathbf{u}+\rho\dot{\mathbf{u}}$, we obtain the temporal equation for the density as
+$\mathbf{u}\dot{\rho}=-\nabla\cdot[(p\mathbf{I}+\underline{\underline{\bm{\Theta}}})]-\rho\nabla\cdot[\tau_{1}(\nabla\bm{\zeta}+\nabla\bm{\zeta}^{T})+\tau_{2}(\nabla\cdot\bm{\zeta})\mathbf{I}].$ (75)
+As shown below, Eq. (75) provides a generalized form of Bernoulli’s law, a general definition of the sound speed, and a generalized equation of state (EOS).
+Considering inviscid fluids via setting $\tau_{1}=\tau_{2}\approx 0$, and the steady state, $\dot{\rho}=\mathbf{u}\cdot\nabla\rho$ where we set $\partial_{t}\rho=0$, we rewrite Eq. (75) as
+$\mathbf{u}(\mathbf{u}\cdot\nabla\rho)=-[\nabla\cdot(p\mathbf{I}+\underline{\underline{\bm{\Theta}}})].$ (76)
+By considering a one-dimensional setup and using the relation $\mathrm{d}(\rho u^{2})=u^{2}\mathrm{d}\rho+2\rho u\mathrm{d}u$, Eq. (76) is further simplified as
+$\mathrm{d}(p+\rho u^{2})=2\rho u\mathrm{d}u.$ (77)
+This result is nothing but a generalized Bernoulli’s equation. When there is no change of the velocity entropy, namely, $\rho u\mathrm{d}u=0$, we have
+$p+\rho u^{2}=\text{constant},$ (78)
+which is Bernoulli’s principle.
+However, when velocity entropy increases, $\mathrm{d}u\neq 0$, the associated energy is changed as well; in this case, the classic Bernoulli’s equation loses its validity.
+VI.4 The problem of the conventional kinetic energy formulation
+Conventional formulation for the kinetic energy is $\tilde{\mathcal{K}}=\int_{\Omega}\rho\mathbf{u}^{2}\mathrm{d}\Omega$. In this case, the time derivative of the kinetic energy reads
+$\frac{\mathrm{d}{\tilde{\mathcal{K}}}}{\mathrm{d}t}=\int_{\Omega}\bigg(\frac{\mathrm{d}\rho}{\mathrm{d}t}\mathbf{u}^{2}+\rho\mathbf{u}\cdot\frac{\mathrm{d}\mathbf{u}}{\mathrm{d}t}\bigg)\mathrm{d}\Omega.$ (79)
 A noteworthy shortcoming of Eq.
@@
 Moreover, the first part in Eq.
-(72) requires a prior knowledge about the time evolution of the density, which is unknown for two immiscible phases.
-The density evolution can be avoided if the kinetic energy is written as $\rho\mathbf{u}\cdot\mathbf{u}=\bm{\zeta}\cdot\mathbf{u}$; in this case, the time evolution of the momentum $\bm{\zeta}$ as a whole term is sufficient, subjected to the momentum conservation.
-Noteworthily, the so-called continuity equation $\partial_{t}\rho+\nabla\cdot(\rho\mathbf{u})=0$ cannot be applied here for density evolution.
-The reason is as follows: (I) The classic continuity equation is derived for pure fluid, where there is no density variation in space.
-(II) When considering two immiscible fluids, the classic continuity equation cannot be straightforwardly adopted, since the density varies across the interface.
-In this case, the mass conservation and energy dissipation are not independent.
-Suffice it to say, one cannot directly apply a mass conservation without considering the energy dissipation across the interface.
-(III) Within the interface of two immiscible fluids, one cannot use the condition $\nabla\cdot\mathbf{u}=0$ which is used for pure fluids.
-The divergence free condition $\nabla\cdot\mathbf{u}=0$ implies a uniform excess volume everywhere in the whole domain $\Omega$ and that the mass is uniformly distributed in the speed of sound.
-In other words, when the fluid velocity is much less than the speed of sound, the density almost has no variation in space; only in this case, the condition $\nabla\cdot\mathbf{u}=0$ can be applied.
-For pure homogeneous fluids, the condition $\nabla\cdot\mathbf{u}=0$ is used to solve the unknown pressure $p_{0}$, which is homogeneous and can be considered as a reference value (Fig.
-3).
-However, for two immiscible fluids, the pressure is inhomogeneous across the interface, which cannot be solved simply by $\nabla\cdot\mathbf{u}=0$ due to the density variation.
-VI.4 Density evolution equation and the generalization of Bernoulli’s law
-Eq. (66) subtracting Eq. (67) with the relation $\dot{\bm{\zeta}}=\dot{\rho}\mathbf{u}+\rho\dot{\mathbf{u}}$, we obtain the temporal equation for the density as
-$\mathbf{u}\dot{\rho}=-\nabla\cdot[(p\mathbf{I}+\underline{\underline{\bm{\Theta}}})]-\rho\nabla\cdot[\tau_{1}(\nabla\bm{\zeta}+\nabla\bm{\zeta}^{T})+\tau_{2}(\nabla\cdot\bm{\zeta})\mathbf{I}].$ (73)
-As demonstrated in the following, Eq. (73) provides a generalization of Bernoulli’s law.
-Considering inviscid fluids via setting $\tau_{1}=\tau_{2}\approx 0$, and the steady state, $\dot{\rho}=\mathbf{u}\cdot\nabla\rho$ where we set $\partial_{t}\rho=0$, we rewrite Eq. (73) as
-$\mathbf{u}^{2}\nabla\rho=-[\nabla\cdot(p\mathbf{I}+\underline{\underline{\bm{\Theta}}})].$ (74)
-By considering a one-dimensional setup and using the relation $\mathrm{d}(\rho u^{2})=u^{2}\mathrm{d}\rho+2\rho udu$, Eq. (74) is further simplified as
-$\mathrm{d}(p+\rho u^{2})=2u\mathrm{d}\rho.$ (75)
-This result is nothing but a generalized Bernoulli’s equation. When there is no density variation, namely, $\mathrm{d}(p+\rho u^{2})=0$, we have
-$p+\rho u^{2}=\text{constant},$ (76)
-which is Bernoulli’s principle. However, when density varies, $\mathrm{d}\rho\neq 0$, the associated energy is changed as well; in this case, the classic Bernoulli’s equation loses its validity.
+(79) requires a prior information about the time evolution of the density, which is unknown for two immiscible phases within the interface, as the excess volume within the interface is unknown.
+It is important to note that the classical continuity equation, $\partial_{t}\rho+\nabla\cdot(\rho\mathbf{u})=0$, cannot be applied here for density evolution.
+The reasons are three-fold:
+(I) The classical continuity equation is derived for a one-component fluid, in which a single variable-either the density $\rho$ or the volume composition $\phi$-is sufficient to characterize the thermodynamic state.
+In such systems, density variations with pressure follow a standard equation of state (EOS), such as the van der Waals EOS.
+However, within the interface between two immiscible fluids, mixing entropy leads to a local excess volume that is generally non-zero and non-uniform in space.
+Under these conditions, the classical van der Waals EOS loses validity because it applies only to ideal mixing, where the excess volume of two mixing components is zero everywhere.
+For non-ideal mixing of two components, three volume contributions must be considered: the component volumes $v_{1}$ and $v_{2}$, and the excess volume $v_{e}$.
+When $v_{e}=0$, the classical EOS may be applied to $v_{1}$ and $v_{2}$, relating the partial pressures $p_{1}$ and $p_{2}$ to the corresponding densities $\rho_{1}$ and $\rho_{2}$.
+When $v_{e}\neq 0$, an additional EOS governing $v_{e}$ is required, and this cannot be handled within the classical framework of continuity equations because an additional degree of freedom due to excess volume has been introduced.
+As discussed in Section III, our approach is to redistribute the excess volume $v_{e}$ into $v_{1}$ and $v_{2}$ and to introduce a new volume concentration that incorporates the excess-volume information.
+With this redistribution, the effective densities $\rho_{1}$ and $\rho_{2}$ are no longer constants; their actual values, including the redistributed excess volume, depend on the chosen redistribution scheme and are not of primary interest here.
+Instead, we derive a new density equation that directly relates density variations to changes in both pressure and velocity entropy [Eq.
+(75)].
+This stands in contrast to the classical continuity equation, in which density variations are independent of pressure and entropy.
+In the classical framework, one must prescribe an EOS externally to obtain the kinetic evolution.
+In the present theory, by contrast, the EOS emerges naturally from the evolution equations themselves (see Section VIB), aided by the pressure obtained from the modified Gibbs–Duhem relation and the newly defined volume concentration that incorporates the excess volume.
+(II) In the case of a spatially uniform density for one-component fluids, the classical continuity equation reduces to $\nabla\cdot\mathbf{u}=0$.
+This divergence-free condition, which characterizes incompressible one-component fluids, does not hold within the interface between two immiscible fluids.
+Enforcing $\nabla\cdot\mathbf{u}=0$ in such regions produces a pressure field that lacks physical meaning.
+When the excess volume is nonzero, the fluid is inherently compressible, and thus deviates from the incompressibility assumption.
+In this case, the pressure associated with the compressibility arising from the excess volume must be determined from the modified Gibbs–Duhem relation, rather than from solving $\nabla\cdot\mathbf{u}=0$.
+As already discussed in Section III, two variables are required to describe the thermodynamic state of the fluids.
+The volume concentration accounts for the compressibility and determines the pressure according to Eq.
+(58).
+The density evolution equation then naturally gives rise to EOS.
+(III) The classical continuity equation cannot incorporate diffusion processes, such as evaporation and condensation, which are incompatible with the incompressibility assumption, as noted in the Introduction and discussed elsewhere Abels et al.
+(2012).
+In fact, the diffusion equation arises from the presence of excess volume or equivalently from the mixing entropy $s_{\phi}$, thereby leading to an effective compressibility.
 VI.5 Replication of the Navier-Stokes equations
@@
 (69) by $\rho$, the dissipation equation reads [a factor $1/2$ has been added due to the kinetic energy formulation, $(1/2)\rho\mathbf{u}\cdot\mathbf{u}$ ]
-$\frac{1}{2}\rho\frac{\mathrm{d}\mathbf{u}}{\mathrm{d}t}=\rho\{\nabla\cdot\tau_{1}[\nabla(\rho\mathbf{u})+\nabla(\rho\mathbf{u})^{T}]+\nabla\cdot\tau_{2}\nabla\cdot(\rho\mathbf{u})\mathbf{I}\}=\nabla\cdot\eta(\nabla\mathbf{u}+\nabla\mathbf{u}^{T})+\nabla\cdot\nu\nabla\cdot\mathbf{u})\mathbf{I}.$ (79)
+$\frac{1}{2}\rho\frac{\mathrm{d}\mathbf{u}}{\mathrm{d}t}=\rho\{\nabla\cdot\tau_{1}[\nabla(\rho\mathbf{u})+\nabla(\rho\mathbf{u})^{T}]+\nabla\cdot\tau_{2}\nabla\cdot(\rho\mathbf{u})\mathbf{I}\}=\nabla\cdot\eta(\nabla\mathbf{u}+\nabla\mathbf{u}^{T})+\nabla\cdot[\nu(\nabla\cdot\mathbf{u})\mathbf{I}].$ (82)
 Summing Eq. (81) with Eq. (82), we obtain the classic Navier-Stokes equation for incompressible fluids as
@@
 We stress that viscous dissipation inherently contradicts the incompressibility condition.
-When viscous dissipation occurs, density cannot remain constant, as dissipative processes—such as heat release, light emission, or electrical discharge—introduce entropy changes that modify the density.
+When viscous dissipation occurs, density cannot remain constant, as dissipative processes introduce entropy changes that modify the density.
 However, under dissipation-free conditions (e.g., isentropic flow), density can reasonably be treated as constant.
@@
 Once the dissipation formulation is known, the time evolution of the density can be deduced.
-The dissipation formulation, Eq. (79) is particularly chosen to be consistent with the Newton viscous stress tensor to dissipate the velocity and the associated energy as
-$\int_{\Omega}\rho\mathbf{u}\cdot\frac{\mathrm{d}\mathbf{u}}{\mathrm{d}t}\mathrm{d}\Omega=-\tau_{1}\int_{\Omega}\nabla(\rho\mathbf{u}):\nabla(\rho\mathbf{u})\mathrm{d}\Omega-\tau_{2}\int_{\Omega}[\nabla\cdot(\rho\mathbf{u})]^{2}\mathrm{d}\Omega\leq 0.$ (81)
-Other dissipation formations for $\mathrm{d}\mathbf{u}/\mathrm{d}t$ are also possible as long as the integration
-$\int_{\Omega}\rho\mathbf{u}\cdot\frac{\mathrm{d}\mathbf{u}}{\mathrm{d}t}\mathrm{d}\Omega$ (82)
-is not greater than 0.
 VI.6 Remarks
-The key point of the present theory is the energy minimization based on an alternative formulation for the kinetic energy, $\rho\mathbf{u}\cdot\mathbf{u}$, rather than the conventional formulation, $\frac{1}{2}\rho\mathbf{u}^{2}$.
+The key point of the present theory is the energy minimization based on an alternative formulation for the kinetic energy, $\rho\mathbf{u}\cdot\mathbf{u}$, rather than the conventional formulation, $\rho\mathbf{u}^{2}$.
 In our derivation, the dissipation of the kinetic energy is divided into two parts.
@@
 This derivation differs from Euler and Navier-Stokes equations where the momentum variation is only caused by the evolution of the fluid velocity.
-The second part of the kinetic energy dissipation is the velocity dissipation due to the difference in the density, viscosity, and velocity, as explained in Fig.
-4III.
-Another vital point of the current work is the pressure energy, the formulation of which is consistent with many pioneers Chen and Doolen (1998).
-The physical meaning of the pressure has been elucidated in Fig.
-3.
-The minimization of the pressure energy leads to the so-called grand potential model Wang et al.
-(2012).
-In the lattice model, the free energy and the pressure can be explicitly formulated according to the theory of thermodynamics, such as Landau potential.
-However, for non-lattice models, the pressure cannot be explicitly identified.
-This case is the consideration in Euler and Navier-Stokes equations, where the unknown pressure is solved by the divergence free of the velocity for incompressible flow.
-For liquid and solid phases, the lattice model can be applied because of the regularized order of the structure within the short range.
-For gas phases, the lattice model loses its validity due to the disorder of the structure; one has to use the equation of state to formulate the partial pressure of the gas phase Wöhrwag et al.
-(2018); Ledesma-Aguilar et al.
-(2014).
+The second part of the kinetic energy dissipation is the velocity dissipation associated with the velocity entropy.
 In the current work, we consider the linear momentum $\bm{\zeta}=\rho\mathbf{u}$.
 Non-linear terms related to fluid velocity may be accounted for the dissipation if we consider non-linear momentum, for example, $\bm{\zeta}\propto\mathbf{u}^{n}$ or $\bm{\zeta}\propto(\nabla\mathbf{u})^{n}$.
-The non-linear momentum not only modifies the velocity dissipation but also changes the momentum balance, which may be applied for understanding highly nonlinear effects, such as turbulence.
+The non-linear momentum not only modifies the velocity dissipation but also changes the momentum balance, which may be applied for understanding highly nonlinear effects.
 These specialized cases are out of the scope of the present discussion.
@@
 (69).
-The local velocity can increase (acceleration) or decrease (deceleration), but overall the kinetic energy $\mathcal{K}$ decreases with time in a closed system.
+The local velocity can increase (acceleration) or decrease (deceleration), but overall the kinetic energy $\mathcal{K}$ plus other potential energies decreases with time in a closed system.
 Once dissipation takes place, i.e., $\mathrm{d}\mathbf{u}/\mathrm{d}t\neq 0$, the density at the interface cannot be assumed to be a constant; in this case, the density evolves with time.
@@
 $\frac{\mathrm{d}\mathcal{K}}{\mathrm{d}t}=\int_{\Omega}\bigg[\frac{\mathrm{d}(\rho\mathbf{u})}{\mathrm{d}t}\cdot\mathbf{u}+\rho\mathbf{u}\cdot\frac{\mathrm{d}\mathbf{u}}{\mathrm{d}t}\bigg]\mathrm{d}\Omega.$ (85)
-The first part corresponds to the momentum balance.
-The second part is associated with the velocity dissipation.
-Overall, $\mathrm{d}\mathcal{K}/\mathrm{d}t\leq 0$, which is the key basis of the present finding.
+The first part corresponds to the momentum balance. The second part is associated with the velocity dissipation.
 VII Case Studies and Applications: Validation and Broader Implications
@@
 which is consistent with the results of the perturbation analysis Landau and Lifshitz (2013).
-Note that two conditions have been applied to derive the sound speed: $\delta q\approx 0$ (see appendix) and $\eta\approx 0$, known as isentropic condition (adiabatic and reversible).
+Note that two conditions have been applied to derive the sound speed: $\delta q\approx 0$ and $\eta\approx 0$, known as isentropic condition (adiabatic and reversible).
 For one component fluid, the Landau potential is expressed as $f-\sum_{i}\mu_{i}\phi_{i}=f$, where the free energy $f$ is written as $f=\theta-Ts$.
@@
 By dividing both sides of Eq.
-(96) by the molar volume, we obtain $\Psi\sim\rho\ln\rho$, as used in Ref.
+(97) by the molar volume, we obtain $\Psi\sim\rho\ln\rho$, which is the prescribed EOS used in Refs.
 Ledesma-Aguilar et al.
-(2014).
+(2014); Ridl and Wagner (2018).
+This highlights our central result: the density evolution in our theory naturally generates the EOS, in contrast to the prior EOS imposed in existing diffuse-interface methods.
 VII.3 Water-air interface
@@
 $\frac{\partial\phi}{\partial t}+\mathbf{u}\cdot\nabla\phi=\nabla\cdot\varsigma\nabla\mu,$ (99)
-$\mu=2\phi(1-\phi)-2\nabla^{2}\phi,$ (99)
+$\mu=2\phi(1-\phi)(1-2\phi)-2\nabla^{2}\phi,$ (100)
 $\mathbf{u}\frac{\mathrm{d}\rho}{\mathrm{d}t}=\mathbf{u}\bigg(\frac{\partial\rho}{\partial t}+\mathbf{u}\cdot\nabla\rho\bigg)=-\nabla p,$ (101)
@@
 $\mu(x)=0,$ (105)
-$u_{0}^{2}\frac{\mathrm{d}\rho}{\mathrm{d}x}=-\frac{\mathrm{d}p}{\mathrm{d}x},$ (105)
+$u^{2}\frac{\mathrm{d}\rho}{\mathrm{d}x}=-\frac{\mathrm{d}p}{\mathrm{d}x},$ (106)
 $u(x)=u_{0}\exp(-ax),$ (107)
@@
 F.W.
-proposed the concept, carried out the theoretical derivations, and prepared the manuscript.
+conceived the overall concept, conducted all theoretical derivations, and prepared the manuscript.
 B.N.
 contributed to securing the funding.
-Appendix A Gibbs-Duhem equation and the kinetic energy
-In the case of incompressible fluids, a single variable—either $\phi$ or $\rho$—is sufficient to characterize the state of the system.
-For compressible fluids with nearly identical densities, a single variable likewise suffices to adequately describe the system, as demonstrated in Section III.
-Here, we illustrate the Gibbs–Duhem equation by taking $\phi$ as the state variable.
-According to the first law of thermodynamics, the change in the internal energy density $\theta$ of a unit volume element $v$ can be expressed as
-$\mathrm{d}\theta=\delta q-p\mathrm{d}v+\sum_{i}\mu_{i}\mathrm{d}\phi_{i}.$ (112)
-The first term corresponds to heat transfer, while the second and third terms represent mechanical and chemical work, respectively.
-Typically, the timescale associated with heat transfer, represented by $\delta q$, is much longer than that of mass transport, as thermal diffusivity exceeds mass diffusivity by at least three orders of magnitude.
-Consequently, we assume an isothermal condition on the timescale of mass transport and adopt $\delta q\approx 0$ in the current work.
-Nonetheless, the temperature evolution equation is derived in Section V to provide a complete description of the system.
-Combining Eq. (112) with the total derivative $\mathrm{d}\theta=-p\mathrm{d}v-v\mathrm{d}p+\sum_{i}\mu_{i}\mathrm{d}\phi_{i}+\sum_{i}\phi_{i}\mathrm{d}\mu_{i}$, we arrive at the Gibbs–Duhem equation:
-$v\mathrm{d}p+\sum_{i}\phi_{i}\mathrm{d}\mu_{i}=0,$ (113)
-which has been often used to the formulate the pressure force for a unit volume
-$\nabla p=\sum_{i}\phi_{i}\nabla\mu_{i}$ (114)
-for the momentum evolution.
-We note that a limitation of the above derivation is the omission of the kinetic energy, $\rho\mathbf{u}^{2}$ due to convection.
-The classical Gibbs–Duhem relation assumes thermodynamic equilibrium at each time step—that is, a quasi-static process—and therefore does not account for the momentum balance associated with kinetic energy.
-Another limitation arises in systems with large density ratios, where two variables, $\rho$ and $\phi$, are required to fully describe the change in system energy.
-In this context, the quintessential Gibbs–Duhem relation considers only the mechanical/chemical work associated with $\mathrm{d}\phi$, while the contribution from $\mathrm{d}\rho$ is neglected.
-A more complete treatment should consider the total energy, including internal energy ($\theta$), kinetic energy ($\rho\mathbf{u}\cdot\mathbf{u}$), and thermodynamic pressure ($p$), i.e., $\mathrm{d}(\theta+p+\rho\mathbf{u}\cdot\mathbf{u})$, rather than only $\mathrm{d}\theta$ as in the quasi-static approximation.
-When the contribution of the mixing entropy $s$ is considered, the internal energy $\theta$ should be replaced by the free energy density $f=\theta-Ts$.
-One illustrative example is that minimizing the sum of the thermodynamic pressure and kinetic energy naturally leads to momentum conservation:
-$\mathrm{d}(p+\rho\mathbf{u}\cdot\mathbf{u})=\mathrm{d}p+\mathrm{d}(\rho\mathbf{u})\cdot\mathbf{u}+\rho\mathbf{u}\cdot\mathrm{d}\mathbf{u}.$ (115)
-In this expression, the first two terms contribute to momentum conservation: $\mathrm{d}(\rho\mathbf{u})/\mathrm{d}t=-\nabla p$ by reformulating $\mathrm{d}p$ as $\mathrm{d}p/\mathrm{d}t=(\mathrm{d}p/\mathrm{d}\mathbf{r})\cdot(\mathrm{d}\mathbf{r}/\mathrm{d}t)=\nabla p\cdot\mathbf{u}$).
-The last term in Eq.
-(115) represents the dissipation of velocity.
-The comprehensive energy formulation encompassing the free energy, pressure energy, kinetic energy, and wall free energy is expressed in Eq.
+Appendix A The modified Gibbs-Duhem equation and the pressure
+The modified Gibbs-Duhem relation reads
+$\delta p+\sum_{i=1}^{K}\phi_{i}\delta\mu_{i}+\rho\mathbf{u}\cdot\delta\mathbf{u}=0.$ (113)
+The velocity term is replaced by the entropy contribution as $\rho\mathbf{u}\cdot\delta\mathbf{u}=-T\delta s_{u}=-\delta(Ts_{u})$.
+In the last equality, we have used the isothermal condition.
+The chemical potential term $\phi_{i}\delta\mu_{i}$ is rewritten as $\phi_{i}\delta\mu_{i}=\delta(\mu_{i}\phi_{i})-\mu_{i}\delta\phi_{i}$ according to the chain rule of total derivative.
+Recalling the definition of the chemical potential, $\mu_{i}=\delta e/\delta\phi_{i}$, we further write the chemical potential term $\phi_{i}\delta\mu_{i}$ as $\sum_{i=1}^{K}\phi_{i}\delta\mu_{i}=\sum_{i=1}^{K}\delta(\mu_{i}\phi_{i})-\delta e$ (see Appendix C).
+Summarizing all these equalities, we obtain the following equation
+$\delta(-p+\sum_{i=1}^{K}\mu_{i}\phi_{i}-e-Ts_{u})=\delta(-p+\sum_{i=1}^{K}\mu_{i}\phi_{i}-e_{0})=0.$ (114)
+In the above equation, we have used the equality $e+Ts_{u}=e_{0}$. Integrating Eq. (114), we obtain the final expression for the pressure as
+$p=e_{0}-\sum_{i=1}^{K}\mu_{i}\phi_{i}+p_{\text{ref}},$ (115)
+where $p_{\text{ref}}$ is a constant. For simplicity, this constant is set as zero in the current work.
+Appendix B $\delta$ operator, $\nabla$ operator, and the formulation of the chemical potential
+If replacing the $\delta$ operator by the space gradient operator $\nabla$ and the velocity term by the velocity entropy $T\nabla s_{u}$ in Eq.
+(113), the following equation for the gradient of the pressure is obtained
+$\nabla p=-\sum_{i=1}^{K}\phi_{i}\nabla\mu_{i}+T\nabla s_{u}.$ (116)
+When using the original definition of the chemical potential as
+$\mu_{i}=\frac{\mathrm{d}e}{\mathrm{d}\phi_{i}}=\partial e/\partial\phi_{i}+(\partial e/\partial\nabla\phi_{i})\cdot\nabla\phi_{i},$ (117)
+which is consistent with Ref. Ridl and Wagner (2018), Eq. (116) is rewritten as
+$\nabla p-\nabla(Ts_{u})+\sum_{i=1}^{K}[\nabla(\mu_{i}\phi_{i})-\mu_{i}\nabla\phi_{i}]=0.$ (118)
+Integrating Eq. (118) by using Eq. (117), we obtain the same formulation of Eq. (115).
+Note that if the conventional functional derivative
+$\mu_{i}=\partial e/\partial\phi_{i}-\kappa_{i}\nabla^{2}\phi_{i}=\partial f_{0}/\partial\phi_{i}-\kappa_{i}\nabla^{2}\phi_{i}$ (119)
+is used Ledesma-Aguilar et al.
+(2014), one cannot obtain Eq.
 (115).
-The detailed derivation of the corresponding energy minimization, which represents a key result of the present work, is presented in Section VI.
+By using Eq.
+(119), one obtains a different formulation of the pressure $p=f_{0}-\sum_{i=1}^{K}\kappa_{i}(\nabla\phi_{i})^{2}-\sum_{i=1}^{K}\mu_{i}\phi_{i}+p_{\text{ref}}$; a negative sign occurs in the second term.
+The reason for this inconsistency is as follows.
+Only after applying the integration by parts to the term $(\partial e/\partial\nabla\phi_{i})\cdot\nabla\phi_{i}$ in Eq.
+(117) on the whole domain $\Omega$, we obtain the second term with a negative sign in Eq.
+(119).
+This negative sign arises from integration by parts and leads to inconsistent results for the pressure.
+One cannot use this result—obtained after applying integration by parts in the whole domain—to integrate the Gibbs–Duhem relation, which is a local differential equation.
+Instead, the original definition given in Eq.
+(117) should be used when integrating the Gibbs–Duhem relation.
+Appendix C Derivative of $\mu\phi$
+Supposing that $\phi(x)$ and $\phi(y)$ are functions: $\mathbb{R}^{3}\mapsto\mathbb{R}$, and $\mu$ is a functional: $\Phi\mapsto\mathbb{R}$, where $\phi(x),\ \phi(y)\in\Phi$, the derivative for the product of $\mu$ and $\phi$ reads Gelfand et al.
+(2000)
+$\frac{\delta[\mu\phi(x)]}{\delta\phi(y)}=\phi(x)\frac{\delta\mu}{\delta\phi(y)}+\mu\frac{\delta\phi(x)}{\delta\phi(y)}.$ (120)
+In a special case, $\phi(x)=\phi(y)$, we have $\frac{\delta\phi(x)}{\delta\phi(y)}=1$, and therefore $\delta(\mu\phi)=\phi\delta\mu+\mu\delta\phi$.
+As $\mu=\delta e/\delta\phi$, the above equation is finally expressed as $\delta(\mu\phi)=\phi\delta\mu+\delta e$.

### v4

@@
     When diffusion takes place, i.e., $\mathrm{d}\phi/\mathrm{d}t\neq 0$, we have $\mathrm{d}\rho/\mathrm{d}t\neq 0$ if $\rho_{w}\neq\rho_{a}$, so that in the linear interpolation approach, the incompressible condition is not compatible with diffusion for two immiscible phases with distinct densities.
-  • [II(B)]: Uniform excess volume. According to the definition of density, we have the following expression for the mixture density at any position $\bm{r}$ in the considered domain $\Omega$ as
+  • [II(B)]: Uniform excess volume.
+According to the definition of density, we have the following expression for the mixture density at any position $\boldsymbol{r}$ in the considered domain $\Omega$ as
     $\rho=\frac{\rho_{w}v_{w}+\rho_{a}v_{a}}{V}=\rho_{w}\frac{v_{w}}{V}+\rho_{a}\frac{v_{a}}{V},$ (6)
@@
 The key issue is that mass conservation has not been connected to the physical principle of dissipation, as will be demonstrated below.
-Furthermore, the mass conservative derivative is problematic for deriving the continuity equation, as illustrated in the following.
+Furthermore, the mass-conservative derivative poses difficulties for deriving the continuity equation, as illustrated below and discussed in detail in Appendix D.
 In the absence of diffusion, the convective mass flux $\mathbf{j}_{c}$ is commonly expressed in the literature as
@@
 Either the variation of $\rho^{*}$ or $\phi$ can then describe the compressibility.
-However, when $\rho_{w}^{\prime}\gg\rho_{a}^{\prime}$, the van der Waals formulation loses its validity and must be generalized to properly describe the evolution of the density field.
+However, when $\rho_{w}^{\prime}\gg\rho_{a}^{\prime}$, the van der Waals formulation loses validity in the absence of a prescribed equation of state (EOS) and must therefore be generalized to accurately describe the evolution of the density field.
 The above concept can be readily extended to the more general case of a multicomponent system, for which the following equality holds:
@@
 For a binary system, we use the notation $\phi_{1}=\phi$, $\phi_{2}=1-\phi$.
-The density of the mixture at any position $\bm{r}\in\Omega$ is denoted by $\rho(\bm{r})$.
+The density of the mixture at any position $\boldsymbol{r}\in\Omega$ is denoted by $\rho(\boldsymbol{r})$.
 IV Dissipation principle
@@
 This is a standard approach in the literature, except for the inclusion of the pressure term, $\mathcal{P}$. The motivation for introducing $\mathcal{P}$ will be presented below.
-The integration over the local free energy density $e(\bm{\phi},\nabla\bm{\phi})$ leads to
-$\mathcal{F}(\bm{\phi},\nabla\bm{\phi})=\int_{\Omega}e(\bm{\phi},\nabla\bm{\phi})\mathrm{d}\Omega.$ (53)
+The integration over the local free energy density $e(\boldsymbol{\phi},\nabla\boldsymbol{\phi})$ leads to
+$\mathcal{F}(\boldsymbol{\phi},\nabla\boldsymbol{\phi})=\int_{\Omega}e(\boldsymbol{\phi},\nabla\boldsymbol{\phi})\mathrm{d}\Omega.$ (53)
 By decomposing the entropy $s$ into thermal entropy $s_{T}$, component-mixing entropy $s_{\phi}$, and velocity entropy $s_{u}$, we write the free energy density as
@@
 The $\mathcal{K}$ term denotes the kinetic energy of the fluid, which is formulated as
-$\mathcal{K}=\int_{\Omega}\bm{\zeta}\cdot\mathbf{u}\mathrm{d}\Omega,$ (55)
-where $\bm{\zeta}$ is the linear momentum density expressed as $\bm{\zeta}=\rho\mathbf{u}$.
+$\mathcal{K}=\int_{\Omega}\boldsymbol{\zeta}\cdot\mathbf{u}\mathrm{d}\Omega,$ (55)
+where $\boldsymbol{\zeta}$ is the linear momentum density expressed as $\boldsymbol{\zeta}=\rho\mathbf{u}$.
 The $\mathcal{P}=\int_{\Omega}p\mathrm{d}\Omega$ term, often overlooked previously and introduced here, describes the pressure energy.
@@
 Next, we evaluate the total time derivative of the free energy $\mathcal{F}$, the pressure energy $\mathcal{P}$, and the kinetic energy $\mathcal{K}$.
-By using the chain rule and noting that the functional $\mathcal{F}$ has two arguments, $\bm{\phi}$ and $\nabla\bm{\phi}$, the total time derivative of the potential free energy $\mathcal{F}$ reads
-$\dot{\mathcal{F}}(\bm{\phi},\nabla\bm{\phi})=\int_{\Omega}\sum_{i=1}^{K}\frac{\partial e}{\partial\phi_{i}}\frac{\mathrm{d}\phi_{i}}{\mathrm{d}t}+\frac{\partial e}{\partial\nabla\phi_{i}}\cdot\frac{\mathrm{d}\nabla\phi_{i}}{\mathrm{d}t}\mathrm{d}\Omega.$ (59)
+By using the chain rule and noting that the functional $\mathcal{F}$ has two arguments, $\boldsymbol{\phi}$ and $\nabla\boldsymbol{\phi}$, the total time derivative of the potential free energy $\mathcal{F}$ reads
+$\dot{\mathcal{F}}(\boldsymbol{\phi},\nabla\boldsymbol{\phi})=\int_{\Omega}\sum_{i=1}^{K}\frac{\partial e}{\partial\phi_{i}}\frac{\mathrm{d}\phi_{i}}{\mathrm{d}t}+\frac{\partial e}{\partial\nabla\phi_{i}}\cdot\frac{\mathrm{d}\nabla\phi_{i}}{\mathrm{d}t}\mathrm{d}\Omega.$ (59)
 By using the vector calculus $\nabla\cdot(\nabla\phi_{i}\otimes\nabla\phi_{i})=\nabla^{2}\phi_{i}\nabla\phi_{i}+(\nabla\phi_{i}\cdot\nabla)\nabla\phi_{i}$, the definition of materials derivative $\dot{\nabla\phi_{i}}=\partial_{t}\nabla\phi_{i}+\mathbf{u}\cdot\nabla\nabla\phi_{i}$, and integration by parts with no-flux boundary condition (more details can be found in Wang et al.
@@
 (2024)), we obtain the following equations
-$\dot{\mathcal{F}}=\int_{\Omega}\sum_{i=1}^{K}\mu_{i}\dot{\phi}_{i}+(\nabla\cdot\underline{\underline{\bm{\Theta}}})\cdot\mathbf{u}\mathrm{d}\Omega;$ (60)
+$\dot{\mathcal{F}}=\int_{\Omega}\sum_{i=1}^{K}\mu_{i}\dot{\phi}_{i}+(\nabla\cdot\underline{\underline{\boldsymbol{\Theta}}})\cdot\mathbf{u}\mathrm{d}\Omega;$ (60)
 $\mu_{i}=\frac{\delta e}{\delta\phi_{i}}=\frac{\partial e}{\partial\phi_{i}}-\nabla\cdot\frac{\partial e}{\partial\nabla\phi_{i}};$ (61)
-$\underline{\underline{\bm{\Theta}}}=\sum_{i}\frac{\partial e}{\partial\nabla\phi_{i}}\otimes\nabla\phi_{i}.$ (62)
+$\underline{\underline{\boldsymbol{\Theta}}}=\sum_{i}\frac{\partial e}{\partial\nabla\phi_{i}}\otimes\nabla\phi_{i}.$ (62)
 The first part of Eq.
@@
 The second part of Eq.
-(60) defines a stress tensor $\underline{\underline{\bm{\Theta}}}$, which is responsible for the transformation of the potential free energy into the macroscopic kinetic energy Jacqmin (1999).
+(60) defines a stress tensor $\underline{\underline{\boldsymbol{\Theta}}}$, which is responsible for the transformation of the potential free energy into the macroscopic kinetic energy Jacqmin (1999).
 This stress tensor is consistent with the Korteweg stress by evaluating the derivative, $\partial_{\nabla\phi_{i}}e=\kappa_{i}\nabla\phi_{i}$.
@@
 $\dot{\mathcal{P}}=\int_{\Omega}\frac{\mathrm{d}p(\mathbf{r})}{\mathrm{d}t}\mathrm{d}\Omega=\int_{\Omega}\nabla p\cdot\dot{\mathbf{r}}\mathrm{d}\Omega=\int_{\Omega}\nabla\cdot(p\mathbf{I})\cdot\mathbf{u}\mathrm{d}\Omega,$ (63)
-$\dot{\mathcal{K}}=\int_{\Omega}(\dot{\bm{\zeta}}\cdot\mathbf{u}+\bm{\zeta}\cdot\dot{\mathbf{u}})\mathrm{d}\Omega,$ (64)
+$\dot{\mathcal{K}}=\int_{\Omega}(\dot{\boldsymbol{\zeta}}\cdot\mathbf{u}+\boldsymbol{\zeta}\cdot\dot{\mathbf{u}})\mathrm{d}\Omega,$ (64)
 where $\mathbf{I}$ is an identity tensor.
@@
 (64) and collecting all the $\cdot\mathbf{u}$ terms, we obtain the total time derivative of the system energy $\mathcal{E}$ as
-$\dot{\mathcal{E}}=\int_{\Omega}\sum_{i=1}^{K}\underbrace{\mu_{i}\dot{\phi_{i}}}_{\text{I}}+\underbrace{[\nabla\cdot(\underline{\underline{\bm{\Theta}}}+p\mathbf{I})+\dot{\bm{\zeta}}]}_{\text{II}}\cdot\mathbf{u}+\underbrace{\bm{\zeta}\cdot\dot{\mathbf{u}}}_{\text{III}}\mathrm{d}\Omega.$ (65)
+$\dot{\mathcal{E}}=\int_{\Omega}\sum_{i=1}^{K}\underbrace{\mu_{i}\dot{\phi_{i}}}_{\text{I}}+\underbrace{[\nabla\cdot(\underline{\underline{\boldsymbol{\Theta}}}+p\mathbf{I})+\dot{\boldsymbol{\zeta}}]}_{\text{II}}\cdot\mathbf{u}+\underbrace{\boldsymbol{\zeta}\cdot\dot{\mathbf{u}}}_{\text{III}}\mathrm{d}\Omega.$ (65)
 Note that, at isothermal condition, the total energy is conserved when defining $\Lambda=\mathcal{E}+Ts_{\phi}+Ts_{u}$, where the entropy energies of component and velocity have been added.
 Within this framework, we write the time derivative of the total energy as
-$\frac{\mathrm{d}\Lambda}{\mathrm{d}t}=\int_{\Omega}\sum_{i=1}^{K}\bigg(\mu_{i}\frac{\mathrm{d}\phi_{i}}{\mathrm{d}t}+T\frac{\mathrm{d}s_{\phi_{i}}}{\mathrm{d}t}\bigg)+[\nabla\cdot(\underline{\underline{\bm{\Theta}}}+p\mathbf{I})+\dot{\bm{\zeta}}]\cdot\mathbf{u}+\bigg(\bm{\zeta}\cdot\frac{\mathrm{d}\mathbf{u}}{\mathrm{d}t}+T\frac{\mathrm{d}s_{u}}{\mathrm{d}t}\bigg)\mathrm{d}\Omega=0.$ (66)
+$\frac{\mathrm{d}\Lambda}{\mathrm{d}t}=\int_{\Omega}\sum_{i=1}^{K}\bigg(\mu_{i}\frac{\mathrm{d}\phi_{i}}{\mathrm{d}t}+T\frac{\mathrm{d}s_{\phi_{i}}}{\mathrm{d}t}\bigg)+[\nabla\cdot(\underline{\underline{\boldsymbol{\Theta}}}+p\mathbf{I})+\dot{\boldsymbol{\zeta}}]\cdot\mathbf{u}+\bigg(\boldsymbol{\zeta}\cdot\frac{\mathrm{d}\mathbf{u}}{\mathrm{d}t}+T\frac{\mathrm{d}s_{u}}{\mathrm{d}t}\bigg)\mathrm{d}\Omega=0.$ (66)
 VI.2 Kinetic equations
@@
 The second part in Eq. (65), II, reveals the momentum conservation
-$\dot{\bm{\zeta}}=\frac{\mathrm{d}(\rho\mathbf{u})}{\mathrm{d}t}=-\nabla\cdot(\underline{\underline{\bm{\Theta}}}+p\mathbf{I}).$ (68)
+$\dot{\boldsymbol{\zeta}}=\frac{\mathrm{d}(\rho\mathbf{u})}{\mathrm{d}t}=-\nabla\cdot(\underline{\underline{\boldsymbol{\Theta}}}+p\mathbf{I}).$ (68)
 The conservative force at the right hand side is responsible for the exchange of the potential energy $\mathcal{F}+\mathcal{P}$ with the kinetic energy $\mathcal{K}$.
 This result is nothing but momentum balance.
-Here, a more clear physical meaning of the force is presented; the force term consists of the pressure $p$ and the stress tensor $\underline{\underline{\bm{\Theta}}}$.
+Here, a more clear physical meaning of the force is presented; the force term consists of the pressure $p$ and the stress tensor $\underline{\underline{\boldsymbol{\Theta}}}$.
 The stress tensor is due to the gradient term in the potential energy $\mathcal{F}$, while the pressure is contributed by the associated free energy.
@@
 (58)].
-The only term left in the time evolution of system free energy, III, is $\int_{\Omega}\bm{\zeta}\cdot\dot{\mathbf{u}}\mathrm{d}\Omega$.
+The only term left in the time evolution of system free energy, III, is $\int_{\Omega}\boldsymbol{\zeta}\cdot\dot{\mathbf{u}}\mathrm{d}\Omega$.
 By applying the same dissipation principle as done for the scalar field in Eq.
 (67), we obtain the dissipation equation for the velocity as
-$\dot{\mathbf{u}}=\nabla\cdot\tau_{1}(\nabla\bm{\zeta}+\nabla\bm{\zeta}^{T})+\nabla\cdot\tau_{2}(\nabla\cdot\bm{\zeta})\mathbf{I}.$ (69)
+$\dot{\mathbf{u}}=\nabla\cdot\tau_{1}(\nabla\boldsymbol{\zeta}+\nabla\boldsymbol{\zeta}^{T})+\nabla\cdot\tau_{2}(\nabla\cdot\boldsymbol{\zeta})\mathbf{I}.$ (69)
 Unlike the dissipation rule for the scalar variable $\phi_{i}$ with the associated potential $\nabla\mu_{i}$ [see Eq.
-(67)], three terms $\nabla\bm{\zeta}$, $\nabla\bm{\zeta}^{T}$, and $(\nabla\cdot\bm{\zeta})\mathbf{I}$ have to be taken into account for the dissipation of a vector $\mathbf{u}$.
+(67)], three terms $\nabla\boldsymbol{\zeta}$, $\nabla\boldsymbol{\zeta}^{T}$, and $(\nabla\cdot\boldsymbol{\zeta})\mathbf{I}$ have to be taken into account for the dissipation of a vector $\mathbf{u}$.
 The physical meaning of the velocity dissipation is the increase of the velocity entropy $s_{u}$.
@@
 (65) can be interpreted as
-$\bm{\zeta}\cdot\frac{\mathrm{d}\mathbf{u}}{\mathrm{d}t}=-T\frac{\mathrm{d}s_{u}}{\mathrm{d}t}.$ (70)
+$\boldsymbol{\zeta}\cdot\frac{\mathrm{d}\mathbf{u}}{\mathrm{d}t}=-T\frac{\mathrm{d}s_{u}}{\mathrm{d}t}.$ (70)
 The left hand side is responsible for the minimization of the kinetic energy, which is equivalent to the right hand side for the maximization of the velocity entropy.
@@
 The dissipation equation, Eq. (69), is essentially a diffusion equation applied to the vector $\mathbf{u}$, as evident from the following comparison:
-$\nabla\cdot\tau_{1}\nabla\bm{\zeta}=\nabla\cdot\tau_{1}\nabla(\partial k/\partial\mathbf{u});\ \bm{\zeta}=\partial k/\partial\mathbf{u};$ (72)
+$\nabla\cdot\tau_{1}\nabla\boldsymbol{\zeta}=\nabla\cdot\tau_{1}\nabla(\partial k/\partial\mathbf{u});\ \boldsymbol{\zeta}=\partial k/\partial\mathbf{u};$ (72)
 $\nabla\cdot\varsigma\nabla\mu=\nabla\cdot\varsigma\nabla(\partial e/\partial\mathbf{\phi});\ \mu=\partial e/\partial\phi.$ (73)
-The linear momentum, $\bm{\zeta}$, plays a role analogous to that of the chemical potential, $\mu$.
+The linear momentum, $\boldsymbol{\zeta}$, plays a role analogous to that of the chemical potential, $\mu$.
 A further question is that if the velocity is a conserved variable or not. If not, the dissipation has to be modified by the non-conserved form as
-$\frac{\mathrm{d}\mathbf{u}}{\mathrm{d}t}=-\tau_{0}\bm{\zeta},$ (74)
+$\frac{\mathrm{d}\mathbf{u}}{\mathrm{d}t}=-\tau_{0}\boldsymbol{\zeta},$ (74)
 The non-conserved form is known as the Allen-Cahn type dissipation following the gradient descent path, being consistent with the Langevin equation.
 VI.3 Density evolution equation and the generalization of Bernoulli’s law
-Eq. (68) subtracting Eq. (69) with the relation $\dot{\bm{\zeta}}=\dot{\rho}\mathbf{u}+\rho\dot{\mathbf{u}}$, we obtain the temporal equation for the density as
-$\mathbf{u}\dot{\rho}=-\nabla\cdot[(p\mathbf{I}+\underline{\underline{\bm{\Theta}}})]-\rho\nabla\cdot[\tau_{1}(\nabla\bm{\zeta}+\nabla\bm{\zeta}^{T})+\tau_{2}(\nabla\cdot\bm{\zeta})\mathbf{I}].$ (75)
+Eq. (68) subtracting Eq. (69) with the relation $\dot{\boldsymbol{\zeta}}=\dot{\rho}\mathbf{u}+\rho\dot{\mathbf{u}}$, we obtain the temporal equation for the density as
+$\mathbf{u}\dot{\rho}=-\nabla\cdot[(p\mathbf{I}+\underline{\underline{\boldsymbol{\Theta}}})]-\rho\nabla\cdot[\tau_{1}(\nabla\boldsymbol{\zeta}+\nabla\boldsymbol{\zeta}^{T})+\tau_{2}(\nabla\cdot\boldsymbol{\zeta})\mathbf{I}].$ (75)
 As shown below, Eq. (75) provides a generalized form of Bernoulli’s law, a general definition of the sound speed, and a generalized equation of state (EOS).
 Considering inviscid fluids via setting $\tau_{1}=\tau_{2}\approx 0$, and the steady state, $\dot{\rho}=\mathbf{u}\cdot\nabla\rho$ where we set $\partial_{t}\rho=0$, we rewrite Eq. (75) as
-$\mathbf{u}(\mathbf{u}\cdot\nabla\rho)=-[\nabla\cdot(p\mathbf{I}+\underline{\underline{\bm{\Theta}}})].$ (76)
+$\mathbf{u}(\mathbf{u}\cdot\nabla\rho)=-[\nabla\cdot(p\mathbf{I}+\underline{\underline{\boldsymbol{\Theta}}})].$ (76)
 By considering a one-dimensional setup and using the relation $\mathrm{d}(\rho u^{2})=u^{2}\mathrm{d}\rho+2\rho u\mathrm{d}u$, Eq. (76) is further simplified as
@@
 (79) requires a prior information about the time evolution of the density, which is unknown for two immiscible phases within the interface, as the excess volume within the interface is unknown.
-It is important to note that the classical continuity equation, $\partial_{t}\rho+\nabla\cdot(\rho\mathbf{u})=0$, cannot be applied here for density evolution.
-The reasons are three-fold:
-(I) The classical continuity equation is derived for a one-component fluid, in which a single variable-either the density $\rho$ or the volume composition $\phi$-is sufficient to characterize the thermodynamic state.
-In such systems, density variations with pressure follow a standard equation of state (EOS), such as the van der Waals EOS.
-However, within the interface between two immiscible fluids, mixing entropy leads to a local excess volume that is generally non-zero and non-uniform in space.
-Under these conditions, the classical van der Waals EOS loses validity because it applies only to ideal mixing, where the excess volume of two mixing components is zero everywhere.
-For non-ideal mixing of two components, three volume contributions must be considered: the component volumes $v_{1}$ and $v_{2}$, and the excess volume $v_{e}$.
-When $v_{e}=0$, the classical EOS may be applied to $v_{1}$ and $v_{2}$, relating the partial pressures $p_{1}$ and $p_{2}$ to the corresponding densities $\rho_{1}$ and $\rho_{2}$.
-When $v_{e}\neq 0$, an additional EOS governing $v_{e}$ is required, and this cannot be handled within the classical framework of continuity equations because an additional degree of freedom due to excess volume has been introduced.
-As discussed in Section III, our approach is to redistribute the excess volume $v_{e}$ into $v_{1}$ and $v_{2}$ and to introduce a new volume concentration that incorporates the excess-volume information.
-With this redistribution, the effective densities $\rho_{1}$ and $\rho_{2}$ are no longer constants; their actual values, including the redistributed excess volume, depend on the chosen redistribution scheme and are not of primary interest here.
-Instead, we derive a new density equation that directly relates density variations to changes in both pressure and velocity entropy [Eq.
-(75)].
-This stands in contrast to the classical continuity equation, in which density variations are independent of pressure and entropy.
-In the classical framework, one must prescribe an EOS externally to obtain the kinetic evolution.
-In the present theory, by contrast, the EOS emerges naturally from the evolution equations themselves (see Section VIB), aided by the pressure obtained from the modified Gibbs–Duhem relation and the newly defined volume concentration that incorporates the excess volume.
-(II) In the case of a spatially uniform density for one-component fluids, the classical continuity equation reduces to $\nabla\cdot\mathbf{u}=0$.
-This divergence-free condition, which characterizes incompressible one-component fluids, does not hold within the interface between two immiscible fluids.
-Enforcing $\nabla\cdot\mathbf{u}=0$ in such regions produces a pressure field that lacks physical meaning.
-When the excess volume is nonzero, the fluid is inherently compressible, and thus deviates from the incompressibility assumption.
-In this case, the pressure associated with the compressibility arising from the excess volume must be determined from the modified Gibbs–Duhem relation, rather than from solving $\nabla\cdot\mathbf{u}=0$.
-As already discussed in Section III, two variables are required to describe the thermodynamic state of the fluids.
-The volume concentration accounts for the compressibility and determines the pressure according to Eq.
-(58).
-The density evolution equation then naturally gives rise to EOS.
-(III) The classical continuity equation cannot incorporate diffusion processes, such as evaporation and condensation, which are incompatible with the incompressibility assumption, as noted in the Introduction and discussed elsewhere Abels et al.
-(2012).
-In fact, the diffusion equation arises from the presence of excess volume or equivalently from the mixing entropy $s_{\phi}$, thereby leading to an effective compressibility.
+It is important to note that the classical continuity equation, $\partial_{t}\rho+\nabla\cdot(\rho\mathbf{u})=0$, cannot be applied here for density evolution (see Appendix D).
 VI.5 Replication of the Navier-Stokes equations
@@
 The second part of the kinetic energy dissipation is the velocity dissipation associated with the velocity entropy.
-In the current work, we consider the linear momentum $\bm{\zeta}=\rho\mathbf{u}$.
-Non-linear terms related to fluid velocity may be accounted for the dissipation if we consider non-linear momentum, for example, $\bm{\zeta}\propto\mathbf{u}^{n}$ or $\bm{\zeta}\propto(\nabla\mathbf{u})^{n}$.
+In the current work, we consider the linear momentum $\boldsymbol{\zeta}=\rho\mathbf{u}$.
+Non-linear terms related to fluid velocity may be accounted for the dissipation if we consider non-linear momentum, for example, $\boldsymbol{\zeta}\propto\mathbf{u}^{n}$ or $\boldsymbol{\zeta}\propto(\nabla\mathbf{u})^{n}$.
 The non-linear momentum not only modifies the velocity dissipation but also changes the momentum balance, which may be applied for understanding highly nonlinear effects.
@@
 (2014); Ridl and Wagner (2018).
-This highlights our central result: the density evolution in our theory naturally generates the EOS, in contrast to the prior EOS imposed in existing diffuse-interface methods.
+This highlights our central result: the density evolution in our theory naturally generates EOS, in contrast to a prescribed EOS in existing diffuse-interface methods.
 VII.3 Water-air interface
@@
 A noteworthy outcome of this work is the derivation of a novel density evolution equation
-$\mathbf{u}\dot{\rho}=-\nabla\cdot[(p\mathbf{I}+\underline{\underline{\bm{\Theta}}})]-\rho\nabla\cdot[\tau_{1}(\nabla\bm{\zeta}+\nabla\bm{\zeta}^{T})+\tau_{2}(\nabla\cdot\bm{\zeta})\mathbf{I}]$ (112)
+$\mathbf{u}\dot{\rho}=-\nabla\cdot[(p\mathbf{I}+\underline{\underline{\boldsymbol{\Theta}}})]-\rho\nabla\cdot[\tau_{1}(\nabla\boldsymbol{\zeta}+\nabla\boldsymbol{\zeta}^{T})+\tau_{2}(\nabla\cdot\boldsymbol{\zeta})\mathbf{I}]$ (112)
 for the high density ratio system, such as water-air. From this equation, we see that the density evolution or the local excess volume is affected by two factors:
-  • the local pressure $p$ and stress tensor $\underline{\underline{\bm{\Theta}}}$;
+  • the local pressure $p$ and stress tensor $\underline{\underline{\boldsymbol{\Theta}}}$;
   • the viscous dissipation.
@@
 Acknowledgements.
-This research is supported by VirtMat project of the Helmholtz association, as part of the programme “MSE-materials science and engineering” number-43.31.01.
-F.W.
-conceived the overall concept, conducted all theoretical derivations, and prepared the manuscript.
-B.N.
-contributed to securing the funding.
+We express our sincere gratitude to the Editor, Dr. Cottin-Bizonne, for facilitating constructive exchanges with the referees, and to all referees for their careful and insightful comments.
 Appendix A The modified Gibbs-Duhem equation and the pressure
@@
 Appendix C Derivative of $\mu\phi$
-Supposing that $\phi(x)$ and $\phi(y)$ are functions: $\mathbb{R}^{3}\mapsto\mathbb{R}$, and $\mu$ is a functional: $\Phi\mapsto\mathbb{R}$, where $\phi(x),\ \phi(y)\in\Phi$, the derivative for the product of $\mu$ and $\phi$ reads Gelfand et al.
+Supposing that $\phi(\mathbf{x})$ and $\phi(\mathbf{y})$ are functions: $\mathbb{R}^{3}\mapsto\mathbb{R}$, and $\mu$ is a functional: $\Phi\mapsto\mathbb{R}$, where $\phi(\mathbf{x}),\ \phi(\mathbf{y})\in\Phi$, the derivative for the product of $\mu$ and $\phi$ reads Gelfand et al.
 (2000)
-$\frac{\delta[\mu\phi(x)]}{\delta\phi(y)}=\phi(x)\frac{\delta\mu}{\delta\phi(y)}+\mu\frac{\delta\phi(x)}{\delta\phi(y)}.$ (120)
-In a special case, $\phi(x)=\phi(y)$, we have $\frac{\delta\phi(x)}{\delta\phi(y)}=1$, and therefore $\delta(\mu\phi)=\phi\delta\mu+\mu\delta\phi$.
+$\frac{\delta[\mu\phi(\mathbf{x})]}{\delta\phi(\mathbf{y})}=\phi(\mathbf{x})\frac{\delta\mu}{\delta\phi(\mathbf{y})}+\mu\frac{\delta\phi(\mathbf{x})}{\delta\phi(\mathbf{y})}.$ (120)
+In a special case, $\phi(\mathbf{x})=\phi(\mathbf{y})$, we have $\frac{\delta\phi(\mathbf{x})}{\delta\phi(\mathbf{y})}=1$, and therefore $\delta(\mu\phi)=\phi\delta\mu+\mu\delta\phi$.
 As $\mu=\delta e/\delta\phi$, the above equation is finally expressed as $\delta(\mu\phi)=\phi\delta\mu+\delta e$.
+Appendix D The problem of the classic continuity equation for compressible fluids
+For one-component fluids, the mass conservation reads
+$\frac{\partial}{\partial t}\int_{V}nm_{0}\mathrm{d}V=-\int_{\Gamma}\mathbf{j}\cdot\mathbf{n}\mathrm{d}\Gamma,$ (121)
+where $n$ is the number of the mass particles, such as molecules or atoms, and $m_{0}$ is the mass per particle.
+The flux $\mathbf{j}$ represents the flow of mass particles across the boundary $\Gamma$, with normal vector $\mathbf{n}$.
+By applying Gauss’s divergence theorem, we obtain
+$\frac{\partial(nm_{0})}{\partial t}+\nabla\cdot\mathbf{j}=0.$ (122)
+To convert Eq. (122) into an evolution equation for the density, both sides must be divided by the volume element $V$. Based on the following calculation for $V$:
+$V=v_{m}+v_{e},$ (123)
+where $v_{m}$ is the volume of molecules or atoms, and $v_{e}$ is the excess volume which can also be interpreted as the vacancy volume, we classify the discussion into two cases: (I) When the vacancy volume is much smaller than the molecular volume, i.e.
+i.e., $v_{e}\ll v_{m}$, as in a one-component solid or a solid-like liquid, the system corresponds to a nearly incompressible phase.
+(II) When the vacancy volume is comparable to the molecular volume, as in a one-component gas, the system corresponds to a compressible fluid.
+In case (I), the system cannot be significantly compressed, and the local volume element $V$ can be considered nearly constant.
+By defining the flux $\mathbf{j}=m_{0}n\mathbf{u}$ and multiplying both sides by the approximately constant factor $1/V\approx 1/v_{m}$, we recover the classic continuity equation
+$\frac{\partial\rho}{\partial t}+\nabla\cdot(\rho\mathbf{u})=0.$ (124)
+In case (II), the local volume element $V$ can vary with time and space. Consequently, the continuity equation, Eq. (124) takes the form:
+$(nm_{0})\frac{\partial(1/V)}{\partial t}+\frac{1}{V}\frac{\partial(nm_{0})}{\partial t}+\nabla\cdot(\frac{nm_{0}}{V}\mathbf{u})=0.$ (125)
+Due to the large amount of vacancies in case II, $(1/V)$ is not guaranteed to be a conserved variable, since vacancies can appear or disappear locally under the influence of pressure.
+As such, conventional models require a prescribed equation of state (EOS) to define $1/V$, for example, $pV=\text{constant}$ for an ideal gas.
+Moreover, when $(1/V)$ is not conserved, the convection flux cannot be formulated as $\mathbf{j}=\rho\mathbf{u}$, because the density $\rho$ contains information from $1/V$.
+Incorporating a non-conserved variable into the mass conservation equation is therefore not appropriate.
+The same issue arises for multicomponent compressible fluids with $K\in\mathbb{N}$ components, where $(1/V)$ is also not conserved.
+The only difference is that the volume element is rewritten as $V=\sum_{i=1}^{K}v_{i}+v_{e}$, where $v_{i}$ is the volume of the $i$-th component.
+In the static state with $\mathbf{u}=\mathbf{0}$, Eq.
+(124) or Eq.
+(125) returns to a trivial equality: $0=0$, which has no physical meaning.
+In this case, we have to consider the microscopic velocity to derive EOS (Section VII B).
+Within a moving interface indicating that $1/V$ is non-uniform in space, the continuity equation Eq.
+(124) cannot be applied, as $\partial_{t}(1/V)+\nabla\cdot(\mathbf{u}/V)$ is not necessary to be zero, since the vacancy can vanish or appear locally in the moving process of the interface.
+In our model, we consider two distinct evolution equations.
+The first is for the volume concentration $\phi$ including the excess volume and the volumes of all other mass species.
+The evolution of $1/V$ is driven by the maximization of the mixing entropy $s_{\phi}$, resulting in a diffusion equation.
+For single-component fluids, this diffusion occurs via the exchange of molecules with vacancies, such as in self-diffusion.
+For multicomponent fluids, diffusion occurs through the exchange of molecules with both vacancies and other molecular species.
+The second evolution equation governs the density.
+For the approximation of inviscid fluids, the density evolution in our model based on Eq.
+(75) is expressed as
+$\mathbf{u}\frac{\mathrm{d}\rho}{\mathrm{d}t}=\mathbf{u}\bigg(\frac{\partial\rho}{\partial t}+\mathbf{u}\cdot\nabla\rho\bigg)=-\nabla p.$ (126)
+A heuristic understanding for Eq.
+(126) is that $\nabla p\propto\nabla(1/V)$ under a prescribed EOS, which accounts for local compression effects on the volume element and thereby extends the classical continuity equation.
+For incompressible fluids, the pressure work is nearly zero, $pdV\approx 0$, and therefore $\nabla p\approx 0$ in the Gibbs-Duhem relation.
+In this case, Eq.
+(126) reduces to the classic continuity equation for incompressible flow.
+This scenario corresponds to $v_{e}=0$ i.e., a rigorously incompressible fluid with no vacancies.
+For multicomponent incompressible fluids, the density evolution of each component $\rho_{i}$ can be solved separately.
+For compressible fluids ($v_{e}\neq 0$), the density evolution of individual components cannot generally be solved separately, since the effect of $v_{e}$ on $v_{1}$, $v_{2}$, $\cdots$, $v_{K}$ is unknown.
+In our model for compressible fluids, Eq.
+(126) naturally gives rise to the equation of state (Section VII B), so a prescribed EOS is not required.
+However, if a prescribed EOS is available Ridl and Wagner (2018), which implicitly prescribes the effect of $v_{e}$, then the density evolution of different components can, of course, be solved separately.